### v2

# On generating mapping class groups by pseudo-Anosov elements

Abstract.
Wajnryb proved that the mapping class group of a closed oriented surface is generated by two elements. We prove that the mapping class group is generated by two pseudo-Anosov elements. In particular, if the genus is greater than or equal to nine, we can take the generators to two conjugate pseudo-Anosov elements with arbitrarily large dilatations. Another result we prove is that the mapping class group is generated by two conjugate reducible but not periodic elements if the genus is greater than or equal to eight. We also give similar results to the first and third results for the hyperelliptic mapping class group when the genus is greater than or equal to one.
1. Introduction
Let $\Sigma_{g}$ be a closed oriented surface of genus $g$, and let $\mathcal{M}_{g}$ be the mapping class group of $\Sigma_{g}$, i.e., the group of isotopy classes of orientation-preserving homeomorphisms of $\Sigma_{g}$. The hyperelliptic mapping class group $\mathcal{H}_{g}$ is the subgroup of $\mathcal{M}_{g}$ consisting of isotopy classes of orientation preserving homeomorphisms on $\Sigma_{g}$ that commute with some fixed hyperelliptic involution $\iota:\Sigma_{g}\to\Sigma_{g}$. We note that $\mathcal{M}_{g}=\mathcal{H}_{g}$ for $g=0,1,2$ and in particular that $\mathcal{M}_{0}=\mathcal{H}_{0}=\{1\}$.
The problem of finding generating sets for groups is a classical one, and has been studied by many authors. The first finite generating set for $\mathcal{M}_{g}$ was given by Dehn [4], and the generating set consists of $2g(g-1)$ Dehn twists. After that, Lickorish [14] showed that $13g-$ 1 Dehn twists generate $\mathcal{M}_{g}$. Moreover, Humphries [7] proved that $12g+$ 1 Dehn twists suffice to generate $\mathcal{M}_{g}$ and that $2g$ (or less) Dehn twists cannot generate $\mathcal{M}_{g}$. It is well-known that $\mathcal{H}_{g}$ is generated by $12g+$ 1 Dehn twists (see [2]). If we do not require the generators of $\mathcal{M}_{g}$ to be Dehn twists then the number of generators can be reduced. In fact, Lu [15] showed that $\mathcal{M}_{g}$ is generated by three elements. Here, since $\mathcal{M}_{g}$ is not cyclic, at least two elements are needed to generate $\mathcal{M}_{g}$. For this, Wajnryb [27] gave a generating set for $\mathcal{M}_{g}$ consisting of two elements for $g\geq 1$, that is, the generating set is minimal. Since then, various minimal generating sets for $\mathcal{M}_{g}$ have been constructed (see, for example, [10, 29, 1]). The problem of finding torsion (or involution) generating sets has also been studied, for example, for finite simple groups. There have been many results on torsion (or involution) generating sets for $\mathcal{M}_{g}$ (see, for example, [17, 20, 16, 3, 9, 10, 19, 5, 12, 11, 28, 29]). In particular, Korkmaz gave a minimal torsion generating set of $\mathcal{M}_{g}$ consisting of two torsion elements of order $24g+$ 2 for $g\geq 3$ [10] (which was improved to order $g$ by Yildiz [29] for $g\geq 6$). A minimal generating set for $\mathcal{H}_{g}$ consisting of two torsion elements of order $24g+$ 2 and $22g+$ 2 was constructed by Korkmaz [10] for $g=1,2$ and Stukow [25] for $g\geq 3$. We note that $\mathcal{H}_{g}$ can not be generated by involutions for all $g\geq 1$ from homological reason by the work of Birman and Hilden [2].
By Nielsen-Thurston classification, every element $f$ in $\mathcal{M}_{g}$ (resp. $\mathcal{H}_{g}$) has a representative homeomorphism $\phi$ which is of one of the following three types: (1) periodic i.e., $\phi^{m}=\mathrm{id}$ for some integer $m>0$, (2) reducible i.e., $\phi$ leaves invariant a finite collection of pairwise disjoint essential simple closed curves on $\Sigma_{g}$, or (3) pseudo-Anosov [6, 26]. An element $f$ in $\mathcal{M}_{g}$ is said to be periodic (resp. reducible or pseudo-Anosov) if a representative of $f$ is periodic (resp. reducible or pseudo-Anosov). Periodic elements in $\mathcal{M}_{g}$ are just torsion elements. Note that there is an element in $\mathcal{M}_{g}$ which is periodic and reducible as shown in Figure 2 in the next section. A simple example of a reducible but not periodic element is a Dehn twist, and a torsion element in $\mathcal{M}_{g}$ of order $24g+$ 2 is periodic but not reducible. On the other hand, if $f$ is a pseudo-Anosov element then it is neither periodic nor reducible.
The minimal torsion generating set in [10] (resp. [29]) consists of two periodic but not reducible elements (resp. two periodic and reducible elements). Our results are motivated by these works. We give minimal generating sets for $\mathcal{H}_{g}$ and $\mathcal{M}_{g}$ consisting of two pseudo-Anosov elements (see Theorem 1.1) and a minimal generating set for $\mathcal{M}_{g}$ consisting of two reducible but not periodic elements (see Theorem 1.3).
Theorem 1.1.
For $g\geq 1$, the hyperelliptic mapping class group $\mathcal{H}_{g}$ and the mapping class group $\mathcal{M}_{g}$ are generated by two pseudo-Anosov elements.
We next focus on minimal normal generating sets for $\mathcal{M}_{g}$. Since any Dehn twists about nonseparating simple closed curves are conjugate, the above-mentioned generating sets given by Lickorish [14] and Humphries [7] (resp. Birman-Hilden [2]) are normal generating sets of $\mathcal{M}_{g}$ (resp. $\mathcal{H}_{g}$). Moreover, the above-mentioned minimal torsion generating set in [10] (resp. [29]) consisting of two periodic but not reducible (resp. periodic and reducible) elements is also a normal one. With this background, Theorem 1.2 (resp. 1.3) below gives a minimal normal generating set for $\mathcal{M}_{g}$ consisting of two pseudo-Anosov (resp. reducible but not periodic) elements. In Theorem 1.4 blow, we give a minimal normal generating set for $\mathcal{H}_{g}$ consisting of two reducible but not periodic elements. However, we do not know whether $\mathcal{H}_{g}$ is generated by two conjugate periodic (resp. pseudo-Anosov) elements.
Another motivation for Theorem 1.2 comes from the work of Lanier and Margalit [13]. To explain this, we introduce some terminologies for pseudo-Anosov elements. Let $g\geq 2$. When $f$ is a pseudo-Anosov element in $\mathcal{M}_{g}$, $f$ has a representative $\phi$ such that there are transverse measured foliations $(\mathcal{F}^{s},\mu_{s})$ and $(\mathcal{F}^{u},\mu_{u})$ on $\Sigma_{g}$, and a real number $\lambda>1$ so that $\phi\cdot(\mathcal{F}^{s},\mu_{s})=(\mathcal{F}^{s},\lambda^{-1}\mu_{s})$ and $\phi\cdot(\mathcal{F}^{u},\mu_{u})=(\mathcal{F}^{u},\lambda\mu_{u})$. Then, $\mathcal{F}^{s}$ and $\mathcal{F}^{u}$ are called the stable and unstable foliations, and $\lambda$ is called the dilatation or stretch factor of $\phi$. Note that two conjugate pseudo-Anosov elements have the same dilatation. With these terminologies, we can state Lanier and Margalit’s result. They showed that $\mathcal{M}_{g}$ is normally generated by pseudo-Anosov elements with arbitrarily large dilatations for each $g\geq 1$ (see, Theorem 1.3 and Proposition 4.2 in [13]). Theorem 1.2 is a refinement of this result in the sense that two such conjugate pseudo-Anosov elements suffice to generate $\mathcal{M}_{g}$ if $g\geq 9$.
Theorem 1.2.
The mapping class group $\mathcal{M}_{g}$ is generated by two conjugate pseudo-Anosov elements with arbitrarily large dilatations if $g\geq 9$, and by three conjugate pseudo-Anosov elements with arbitrarily large dilatations if $g\geq 3$.
Theorem 1.3.
The mapping class group $\mathcal{M}_{g}$ is generated by two conjugate reducible but not periodic elements if $g\geq 8$, and by three conjugate reducible but not periodic elements if $g\geq 3$.
Theorem 1.4.
The hyperelliptic mapping class group $\mathcal{H}_{g}$ is generated by two conjugate reducible but not periodic elements if $g\geq 1$.
It is well-known that the mapping class group $\mathcal{M}_{g}$ of $\Sigma_{g}$ can be identified with the orbifold fundamental group of the moduli space $\mathbb{M}_{g}$ of Riemann surfaces of genus $g$. When we equiv $\mathbb{M}_{g}$ with the Teichmüller metric, pseudo-Anosov elements in $\mathcal{M}_{g}$ correspond to geodesic loops in $\mathbb{M}_{g}$. In particular, the logarithm of the dilatation of a pseudo-Anosov element in $\mathcal{M}_{g}$ is the length of a geodesic loop in $\mathbb{M}_{g}$. Then, we can rewrite Theorem 1.2 in terms of the geometry of normal covers of $\mathbb{M}_{g}$. Theorem 1.2 implies that we can take two (resp. three) geodesic loops in $\mathbb{M}_{g}$ with the same arbitrarily length that generate the orbifold fundamental group of $\mathbb{M}_{g}$ if $g\geq 9$ (resp. $g\geq 3$).
Acknowledgements. The first author was supported by JSPS KAKENHI Grant Numbers JP20K03618. The second author was supported by JSPS KAKENHI Grant Numbers JP20K03613. The authors would like to thank Eiko Kin and Mustafa Korkmaz for their comments on an earlier version of this paper.
2. Preliminaries
In this section, we recall the basic facts about the mapping class group and present examples of pseudo-Anosov elements and reducible but not periodic elements. More details can be found in [6]. Let $\mathcal{M}_{g}$ be the mapping class group of the closed oriented surface $\Sigma_{g}$ of genus $g$. In this paper, we use the same symbol for a diffeomorphism and its isotopy class, or a simple closed curve and its isotopy class. For a simple closed curve $c$ on $\Sigma_{g}$, we denote by $t_{c}$ the right-handed Dehn twist about $c$. For $f_{1}$ and $f_{2}$ in $\mathcal{M}_{g}$, the notation $f_{2}f_{1}$ means that we first apply $f_{1}$ and then $f_{2}$.
Let $a,b,c,d$ be simple closed curves on $\Sigma_{g}$. Throughout this paper, we will use the following four relations in $\mathcal{M}_{g}$ repeatedly.
  • For any $f\in\mathcal{M}_{g}$, we have $t_{f(a)}=ft_{a}f^{-1}$.
  • If $a$ is disjoint from $b$, then $t_{a}t_{b}=t_{b}t_{a}$, $t_{b}(a)=a$ and $t_{a}(b)=b$.
  • If $a$ intersects $b$ transversely at exactly one point, then the braid relation $t_{a}t_{b}t_{a}=t_{b}t_{a}t_{b}$ holds. This gives $t_{b}t_{a}(b)=a$ (from $t_{t_{b}t_{a}(b)}=(t_{b}t_{a})t_{b}(t_{b}t_{a})^{-1}=t_{a}$), and hence $t_{a}(b)=t_{b}^{-1}(a)$.
  • Let $x,y,z$ be the interior curves on a subsurface of genus 00 with four boundary curves $a,b,c,d$ on $\Sigma_{g}$ as in Figure 1. Then, the lantern relation $t_{d}t_{c}t_{b}t_{a}=t_{z}t_{y}t_{x}$ holds. The lantern relation was discovered by Dehn [4] and rediscovered by Johnson [8].
We assume that the surface of this paper is the $yz$-plane and that $\Sigma_{g}$ is embedded in $\mathbb{R}^{3}$ as in Figure 2 (resp. Figure 3) such that it is invariant under the rotation $r$ by $-\frac{2\pi}{g}$ about the $x$-axis (resp. the rotation $\iota$ by $-\pi$ about the $y$-axis, i.e., the hyperelliptic involution, and the rotation $r_{\pi}$ by $-\pi$ about the $x$-axis). Note that $r_{\pi}$ is an element in $\mathcal{H}_{g}$ since $\iota r_{\pi}=r_{\pi}\iota$.
Figure 1. Simple closed curves $a,b,c,d,x,y,z$ on $\Sigma_{g}$ for the lantern relation.
The notations $a_{i},b_{i},c_{i}$ (resp. $\alpha_{j}$, $\beta$) will always denote the simple closed curves on $\Sigma_{g}$ shown in Figure 2 (resp. 3) for $i=1,2,\ldots,g$ (resp. $j=1,2,\ldots,2g+1$). Note that $\alpha_{1}=a_{1}$, $\alpha_{2i}=b_{i}$, $\alpha_{2i+1}=c_{i}$ for $i=1,2,\ldots,g-1$.
Figure 2. The rotation $r:\Sigma_{g}\to\Sigma_{g}$ by $-\frac{2\pi}{g}$ about the $x$-axis and the simple closed curves $a_{i},b_{i},c_{i}$ on $\Sigma_{g}$ for $i=1,2,\ldots,g$.
Figure 3. The hyperelliptic involution $\iota:\Sigma_{g}\to\Sigma_{g}$, the rotation $r_{\pi}:\Sigma_{g}\to\Sigma_{g}$ by $-\pi$ about the $x$-axis, and the simple closed curves $\alpha_{1},\alpha_{2},\ldots,\alpha_{2g+1},\beta$ on $\Sigma_{g}$.
Theorem 2.1 ([14]).
For $g\geq 1$, $\mathcal{M}_{g}$ is generated by $t_{a_{1}},\ldots,t_{a_{g}},t_{b_{1}},\ldots,t_{b_{g}}$ and $t_{c_{1}},\ldots,t_{c_{g-1}}$.
Theorem 2.2 ([7]).
For $g\geq 1$, $\mathcal{M}_{g}$ is generated by $t_{\alpha_{1}},t_{\alpha_{2}},\ldots,t_{\alpha_{2g}}$ and $t_{\beta}$.
Theorem 2.3 below is very useful to construct a pseudo-Anosov element. A multicurves in $\Sigma_{g}$ is the union of a finite collection of disjoint essential simple closed curves in $\Sigma_{g}$.
Theorem 2.3 (Theorem 3.1 in [23]).
Let $\mathcal{C}=\{\gamma_{1},\gamma_{2},\ldots,\gamma_{m}\}$ and $\mathcal{D}=\{\delta_{1},\delta_{2},\ldots,\delta_{n}\}$ be multicurves in $\Sigma_{g}$ such that no two component curves form a bigon and such that $\Sigma_{g}-\mathcal{C}\cup\mathcal{D}$ is a union of disks. Any product of positive powers of the Dehn twist $t_{\gamma_{i}}$ and negative powers of the Dehn twist $t_{\delta_{j}}$, where each $\gamma_{i}$ and each $\delta_{j}$ appear at least once, is pseudo-Anosov.
Lemma 2.4.
We define
$h_{2i-1}=t_{\alpha_{2i-1}}t_{\alpha_{2i}}^{-1}\cdot t_{\alpha_{2i+1}}t_{\alpha_{2i+2}}^{-1}\cdots t_{\alpha_{2g-1}}t_{\alpha_{2g}}^{-1}\cdot t_{\alpha_{2g+1}}\ \ \mathrm{and}$
$h_{2i}=t_{\alpha_{2i}}^{-1}\cdot t_{\alpha_{2i+1}}t_{\alpha_{2i+2}}^{-1}\cdot t_{\alpha_{2i+3}}t_{\alpha_{2i+4}}^{-1}\cdots t_{\alpha_{2g-1}}t_{\alpha_{2g}}^{-1}\cdot t_{\alpha_{2g+1}}$
for $i=1,2,\ldots,g$ and $h_{2g+1}=t_{\alpha_{2g+1}}$. Moreover, we set
$f_{2i-1}=t_{\beta}\cdot t_{\alpha_{2i-1}}t_{\alpha_{2i}}^{-1}\cdot t_{\alpha_{2i+1}}t_{\alpha_{2i+2}}^{-1}\cdots t_{\alpha_{2g-1}}t_{\alpha_{2g}}^{-1}\ \ \mathrm{and}$
$f_{2i}=t_{\beta}t_{\alpha_{2i}}^{-1}\cdot t_{\alpha_{2i+1}}t_{\alpha_{2i+2}}^{-1}\cdot t_{\alpha_{2i+3}}t_{\alpha_{2i+4}}^{-1}\cdots t_{\alpha_{2g-1}}t_{\alpha_{2g}}^{-1}$
for $i=1,2,\ldots,g$. Then, $h_{1}$ and $h_{2}$ are pseudo-Anosov elements in $\mathcal{H}_{g}$, and $f_{1}$ and $f_{2}$ are pseudo-Anosov elements in $\mathcal{M}_{g}$.
Proof.
Since $t_{\alpha_{1}},t_{\alpha_{2}},\ldots,t_{\alpha_{2g+1}}$ are generators of $\mathcal{H}_{g}$ (see [2]), $h_{1}$ is also in $\mathcal{H}_{g}$. In the notation of Theorem 3.1 in [23], let $\mathscr{C}=\{\alpha_{1},\alpha_{3},\ldots,\alpha_{2g-1},\alpha_{2g+1}\}$ and $\mathscr{D}=\{\alpha_{2},\alpha_{4},\ldots,\alpha_{2g}\}$ be two multi-curves on $\Sigma_{g}$ (i.e., a set of pairwise disjoint simple closed curves). Then, these two multi-curves $\mathscr{C}$ and $\mathscr{D}$ satisfy the condition of Theorem 2.3, and therefore, we see that $h_{1}$ is a pseudo-Anosov element. By applying these arguments again, with $\mathscr{C}=\{\alpha_{1},\alpha_{3},\ldots,\alpha_{2g-1},\alpha_{2g+1}\}$ replaced by $\mathscr{C}^{\prime}=\{\alpha_{3},\alpha_{5},\ldots,\alpha_{2g+1}\}$, we see that $h_{2}$ is also a pseudo-Anosov in $\mathcal{H}_{g}$.
Similarly, when we take two multi-curves $\mathscr{C}=\{\alpha_{1},\alpha_{3},\ldots,\alpha_{2g-1},\beta\}$ (resp. $\mathscr{C}=\{\alpha_{3},\alpha_{5},\ldots,\alpha_{2g-1},\beta\}$) and $\mathscr{D}=\{\alpha_{2},\alpha_{4},\ldots,\alpha_{2g}\}$ for $f_{1}$ (resp. $f_{2}$), we see that $f_{1}$ (resp. $f_{2}$) is a pseudo-Anosov in $\mathcal{M}_{g}$. ∎
Throughout this section, we use the notation $\prod_{i=1}^{k}f_{i}=f_{1}f_{2}\cdots f_{k}$ for a product $f_{1}f_{2}\cdots f_{k}$ of $k$ elements $f_{1},f_{2},\ldots,f_{k}$ in $\mathcal{M}_{g}$.
Lemma 2.5.
Set $\rho_{n}=rt_{c_{1}}t_{b_{1}}^{-n}t_{a_{1}}$ for a positive integer $n$. Then, $\rho_{n}$ is pseudo-Anosov in $\mathcal{M}_{g}$ for $g\geq 3$. Furthermore, for any positive real number $\lambda$, there is an integer $n$ such that the dilatation of $\rho_{n}$ is more than $\lambda$.
Proof.
We note that $\rho_{n}^{g}=\left(\prod_{i=1}^{g}r^{i}t_{c_{1}}t_{b_{1}}^{-n}t_{a_{1}}r^{-i}\right)\cdot r^{g}$ and $r^{g}=\mathrm{id}$. This gives $\rho_{n}^{g}=\prod_{i=1}^{g}r^{i}t_{c_{1}}t_{b_{1}}^{-n}t_{a_{1}}r^{-i}$, and hence $\rho_{n}^{g}=\prod_{i=1}^{g}r^{i}t_{c_{1}}r^{-i}\cdot r^{i}t_{b_{1}}^{-n}r^{-i}\cdot r^{i}t_{a_{1}}r^{-i}$. Since $r(a_{i},b_{i},c_{i})=(a_{i+1},b_{i+1},c_{i+1})$ for $i=1,2,\ldots,g-1$ and $r(a_{g},b_{g},c_{g})=(a_{1},b_{1},c_{1})$, we have
$\rho_{n}^{g}=\prod_{i=1}^{g}t_{r^{i}(c_{1})}t_{r^{i}(b_{1})}^{-n}t_{r^{i}(a_{1})}=t_{a_{2}}t_{b_{2}}^{-n}t_{c_{2}}\cdot t_{a_{3}}t_{b_{3}}^{-n}t_{c_{3}}\cdots t_{a_{g}}t_{b_{g}}^{-n}t_{c_{g}}\cdot t_{a_{1}}t_{b_{1}}^{-n}t_{c_{1}}.$
If we take two multicurves $\mathscr{C}=\{a_{1},c_{1},a_{2},c_{2},\ldots,a_{g},c_{g}\}$ and $\mathscr{D}=\{b_{1},b_{2},\ldots,b_{g}\}$ then by Theorem 2.3, we see that $\rho_{n}^{g}$ is a pseudo-Anosov element. Therefore $\rho_{n}$ is a pseudo-Anosov.
In the case where $g=3$, we will show that, for any positive real number $\lambda$, there is an integer $n$ such that the dilatation of $\rho_{n}$ is more than $\lambda$. Let $\tau$ be the train track on $\Sigma_{3}$ which carries each element of $\mathscr{C}\cup\mathscr{D}$. As shown Figure 4, the train track $\tau$ has 27 branches $\delta_{i}$, for $i=1,\ldots,9$, $\mu_{3j-2}^{k}$, for $j=1,2,3,k=1,2,3$, $\mu_{3j-1}$, for $j=1,2,3$ and $\mu_{3j}^{k}$, for $j=1,2,3,k=1,2$, where the transverse measure for each branch is indicated by the same symbol. The train track $\rho_{n}^{3}(\tau)$ is carried by $\tau$, and $t_{a_{i}}t_{b_{i}}^{-n}t_{c_{i}}(\tau)$ are also carried by $\tau$ for $i=1,\ldots,g$. The set $H$ of transverse mesures on $\tau$ determined by the equations $\mu_{3j-2}^{1}=\mu_{3j-2}^{2}=\mu_{3j-2}^{3},\mu_{3j}^{1}=\mu_{3j}^{2}$ is preserved by the action of $t_{a_{i}}$ $t_{b_{i}}^{-1}$ and $t_{c_{i}}$. We denote $\mu_{3j-2}^{1},\mu_{3j-2}^{2},\mu_{3j-2}^{3}$ by $\mu_{3j-2}$ and $\mu_{3j}^{1},\mu_{3j}^{2}$ by $\mu_{3j}$ for short. The actions of $\rho_{n}^{3}$ and $t_{a_{i}}t_{b_{i}}^{-n}t_{c_{i}}$ for $i=1,\ldots,3$ on $H$ are determined by the actions of them on values $\mu_{j}$, for $j=1,\ldots,9$. For the homeomorphism $\varphi$ on $\Sigma_{3}$ such that $\varphi(\tau)$ is carried by $\tau$, let $M_{\varphi}$ be the incident matrix for $\varphi$, that is, the $(i,j)$ entry $(M_{\varphi})_{i,j}$ of $M_{\varphi}$ is the times $\varphi(\eta_{j})$ passing through $\mu_{i}$. Then the matrix $M_{\rho_{n}^{3}}$ is a non-negative irreducible matrix and the maximal eigenvalue $\mathrm{PF}(M_{\rho_{n}^{3}})$ of $M_{\rho_{n}^{3}}$ equals the dilatation $\lambda_{n}^{3}$ of $\rho_{n}^{3}$, where $\lambda_{n}$ is the dilatation of $\rho_{n}$. Let $M$ and $N$ be matrices with the same size. If, for each $(i,j)$ entries $M_{i,j}$ and $N_{i,j}$ of $M$ and $N$ respectively, $M_{i.j}\geq N_{i,j}$, then we write $M\geq N$. If $\psi=t_{a_{i}}$ or $t_{b_{i}}^{-1}$ or $t_{c_{i}}$, then $M_{\psi}=I_{9}+P$, where $I_{9}$ is the identity matrix of order 9 and $P$ is a matrix all of whose entries are non-negative. Therefore,
$M_{\rho_{n}^{3}}\geq M_{t_{b_{1}}^{-n}}M_{t_{c_{1}}}$ $=\left(\begin{array}[]{ccc|ccc|ccc}1&n&n&0&0&0&0&0&n\\ 0&1&0&0&0&0&0&0&0\\ 0&0&1&0&0&0&0&0&0\\ \hline\cr 0&0&0&1&0&0&0&0&0\\ 0&0&0&0&1&0&0&0&0\\ 0&0&0&0&0&1&0&0&0\\ \hline\cr 0&0&0&0&0&0&1&0&0\\ 0&0&0&0&0&0&0&1&0\\ 0&0&0&0&0&0&0&0&1\end{array}\right)\left(\begin{array}[]{ccc|ccc|ccc}1&0&0&0&0&0&0&0&0\\ 0&1&0&0&0&0&0&0&0\\ 1&0&1&1&0&0&0&0&0\\ \hline\cr 0&0&0&1&0&0&0&0&0\\ 0&0&0&0&1&0&0&0&0\\ 0&0&0&0&0&1&0&0&0\\ \hline\cr 0&0&0&0&0&0&1&0&0\\ 0&0&0&0&0&0&0&1&0\\ 0&0&0&0&0&0&0&0&1\end{array}\right)$
$=\left(\begin{array}[]{ccc|ccc|ccc}n+1&n&n&n&0&0&0&0&n\\ 0&1&0&0&0&0&0&0&0\\ 1&0&1&1&0&0&0&0&0\\ \hline\cr 0&0&1&1&0&0&0&0&0\\ 0&0&0&0&1&0&0&0&0\\ 0&0&0&0&0&1&0&0&0\\ \hline\cr 0&0&0&0&0&0&1&0&0\\ 0&0&0&0&0&0&0&1&0\\ 0&0&0&0&0&0&0&0&1\end{array}\right)\geq\left(\begin{array}[]{ccc|ccc|ccc}n+1&0&0&0&0&0&0&0&0\\ 0&1&0&0&0&0&0&0&0\\ 0&0&1&0&0&0&0&0&0\\ \hline\cr 0&0&0&1&0&0&0&0&0\\ 0&0&0&0&1&0&0&0&0\\ 0&0&0&0&0&1&0&0&0\\ \hline\cr 0&0&0&0&0&0&1&0&0\\ 0&0&0&0&0&0&0&1&0\\ 0&0&0&0&0&0&0&0&1\end{array}\right).$
It is well-known that, if $M$ is a non-negative irreducible matrix and $N$ is a non-negative matrix such that $M\geq N$, then every eigenvalue $\sigma$ of $N$ satisfies $\mathrm{PF}(M)\geq|\sigma|$ (see Theorem 9.1.1 of [24], for example). Therefore, $\lambda_{n}^{3}\geq n+1$.
We can show the same result for the case where $g\geq 4$, as above. ∎
Figure 4. The $x$-axis crosses this paper from the back side to this side. LEFT: simple closed curves $a_{i},b_{i},c_{i}$ $(i=1,2,\ldots,g)$ RIGHT: a train track on $\Sigma_{g}$ carries simple closed curves $a_{i},b_{i},c_{i}$ $(i=1,2,\ldots,g)$
Lemma 2.6.
Set $\rho=rt_{a_{1}}$ and $\rho^{\prime}=rt_{c_{1}}$. Then, $\rho$ and $\rho^{\prime}$ are reducible but not periodic in $\mathcal{M}_{g}$.
Proof.
We have
$(\rho)^{g}$ $=(rt_{a_{1}})^{g}=\left(\prod_{i=1}^{g}r^{i}t_{a_{1}}r^{-i}\right)\cdot r^{g}=\left(\prod_{i=1}^{g}t_{r^{i}(a_{1})}\right)=t_{a_{2}}t_{a_{3}}\cdots t_{a_{g}}t_{a_{1}},$
and hence we see that for any integer $n$, $(\rho)^{gn}=t_{a_{1}}^{n}\cdots t_{a_{g}}^{n}$. This means that $\rho$ is a reducible and not a periodic element. Similar arguments apply to $\rho^{\prime}$. ∎
Lemma 2.7.
Set $\mathcal{R}=(t_{\alpha_{1}}^{2}t_{\alpha_{2}}t_{\alpha_{3}}\cdots t_{\alpha_{2g-1}})^{g-1}t_{\alpha_{2g+1}}^{-1}$. Then, $\mathcal{R}$ is reducible but not periodic in $\mathcal{H}_{g}$. In particular, the relation $t_{\alpha_{2g+1}}=\mathcal{R}^{-(2g-1)}$ holds.
The element $\mathcal{R}^{-1}$ is called a root of Dehn twist of degree $12g-$ 1 (see [22, 21, 18]).
Proof.
Note that $\mathcal{R}$ is in $\mathcal{H}_{g}$ since $t_{\alpha_{1}},t_{\alpha_{2}},\ldots,t_{\alpha_{2g+1}}$ are generators of $\mathcal{H}_{g}$. It is well-know that the relation $(t_{\alpha_{1}}^{2}t_{\alpha_{2}}t_{\alpha_{3}}\cdots t_{\alpha_{2g-1}})^{2g-1}=t_{\alpha_{2g+1}}^{2}$ holds in $\mathcal{H}_{g}$ (see Section 4.4.1 in [6]). This relation is called the chain relation. Since $\alpha_{2g+1}$ is disjoint from $\alpha_{1},\ldots,\alpha_{2g-1}$, by the chain relation, we have
$\mathcal{R}^{2g-1}=(t_{\alpha_{1}}^{2}t_{\alpha_{2}}t_{\alpha_{3}}\cdots t_{\alpha_{2g-1}})^{(g-1)(2g-1)}t_{\alpha_{2g+1}}^{-(2g-1)}=t_{\alpha_{2g+1}}^{2(g-1)}t_{\alpha_{2g+1}}^{-(2g-1)}=t_{\alpha_{2g+1}}^{-1}.$
Since a Dehn twist in $\mathcal{H}_{g}$ is reducible but not periodic in $\mathcal{H}_{g}$, $\mathcal{R}$ is also reducible but not periodic. ∎
3. Proofs of Theorems 1.1 and 1.4
In this section, we prove Theorems 1.1 and  1.4 in the introduction.
Recall that the notations $\alpha_{1},\alpha_{2},\ldots,\alpha_{2g+1}$ and $\beta$ denote the simple closed curves on $\Sigma_{g}$ as in Figure 3. Note that $\alpha_{i}$ (resp. $\beta$) intersects $\alpha_{i+1}$ (resp. $\alpha_{4}$) transversely at exactly one point for $i=1,2,\ldots,2g-1$. Moreover, $\alpha_{i}$ (resp. $\beta$) is disjoint from $\alpha_{j}$ if $|i-j|>1$ (resp. $j\neq 4$). In this section, we will use the following relations
$t_{t_{\alpha_{i+1}}^{-1}(\alpha_{i})}=t_{t_{\alpha_{i}}(\alpha_{i+1})}$ $\mathrm{and}$ $t_{t_{\alpha_{i+1}}(\alpha_{i})}=t_{t_{\alpha_{i}}^{-1}(\alpha_{i+1})}.$
Theorem 1.1 follows from Theorems 3.1 and  3.2 below.
Theorem 3.1.
Let $H$ be the subgroup of $\mathcal{H}_{g}$ generated by $h_{1}$ and $h_{2}$ defined in Lemma 2.4. Then, $H=\mathcal{H}_{g}$ for $g\geq 1$.
Proof.
Note that if a Dehn twist $t_{c}$ about a simple closed curve $c$ and an element $f$ are in $H$, then $t_{f}(c)$ is also in $H$ by the relation $t_{f(c)}=ft_{c}f^{-1}$.
We have $t_{\alpha_{1}}=h_{1}h_{2}^{-1}\in H$, and hence $t_{h_{2}(\alpha_{1})}$ is in $H$ by $h_{2}\in H$. Since $h_{2}(\alpha_{1})=t_{\alpha_{2}}^{-1}(\alpha_{1})$ and $t_{\alpha_{2}}^{-1}(\alpha_{1})=t_{\alpha_{1}}(\alpha_{2})$, we obtain
$t_{h_{2}(\alpha_{1})}=t_{t_{\alpha_{2}}^{-1}(\alpha_{1})}=t_{t_{\alpha_{1}}(\alpha_{2})}=t_{\alpha_{1}}t_{\alpha_{2}}t_{\alpha_{1}}^{-1},$
and therefore, $t_{\alpha_{1}}t_{\alpha_{2}}t_{\alpha_{1}}^{-1}$ is in $H$. By $t_{\alpha_{1}}\in H$, we have $t_{\alpha_{2}}\in H$. We set $h_{3}=t_{\alpha_{2}}h_{2}$, then $h_{3}$ is in $H$, and hence $t_{h_{3}(\alpha_{2})}$ is in $H$. Since $h_{3}(\alpha_{2})=t_{\alpha_{3}}(\alpha_{2})$ and $t_{\alpha_{3}}(\alpha_{2})=t_{\alpha_{2}}^{-1}(\alpha_{3})$, we get
$H\ni t_{h_{3}(\alpha_{2})}=t_{t_{\alpha_{3}}(\alpha_{2})}=t_{t_{\alpha_{2}}^{-1}(\alpha_{3})}=t_{\alpha_{2}}^{-1}t_{\alpha_{3}}t_{\alpha_{2}},$
and therefore $t_{\alpha_{2}}^{-1}t_{\alpha_{3}}t_{\alpha_{2}}$ is in $H$. By $t_{\alpha_{2}}\in H$, we have $t_{\alpha_{3}}\in H$. Repeating this argument, we can show that $t_{\alpha_{1}},\ldots,t_{\alpha_{2g+1}}$ are contained in $H$. Since $\mathcal{H}_{g}$ is generated by these Dehn twists, we see that $H=\mathcal{H}_{g}$, and the proof is complete. ∎
Theorem 3.2.
Let $G$ be the subgroup of $\mathcal{M}_{g}$ generated by $f_{1}$ and $f_{2}$ defined in Lemma 2.4. Then, $G=\mathcal{M}_{g}$ for $g\geq 3$.
Lemmas 3.3 and 3.4 below are used to show Theorem 3.2. In the proofs of Lemmas 3.3 and 3.4 and Theorem 3.2, we will repeatedly use the fact that if a Dehn twist $t_{c}$ about a simple closed curve $c$ and an element $f$ are in $G$, then $t_{f}(c)$ is also in $G$ by the relation $t_{f(c)}=ft_{c}f^{-1}$.
Lemma 3.3.
For $g\geq 3$, the elements $t_{\alpha_{1}},t_{\alpha_{2}},t_{\alpha_{3}},t_{t_{\alpha_{4}}(\alpha_{5})},t_{t_{\alpha_{5}}(\alpha_{6})},f_{3},f_{4}$ and $f_{5}$ are contained in $G$.
Proof.
Note that we have $t_{\alpha_{1}}=t_{\beta}t_{\alpha_{1}}t_{\beta}^{-1}=f_{1}f_{2}^{-1}$ (since $\alpha_{1}$ is disjoint from $\beta$). This means that $t_{\alpha_{1}}$ is in $G$, and hance $t_{\alpha_{1}}^{-1}f_{2}$ is also in $G$. Since we find that $f_{2}(\alpha_{1})=t_{\alpha_{2}}^{-1}(\alpha_{1})=t_{\alpha_{1}}(\alpha_{2})$, we have $t_{\alpha_{1}}^{-1}f_{2}(\alpha_{1})=\alpha_{2}$. Therefore, we get
$t_{\alpha_{2}}\in G$
by $t_{\alpha_{1}},t_{\alpha_{1}}^{-1}f_{2}\in G$. Moreover, by $f_{3}=t_{\alpha_{2}}f_{2}$ (from the fact that $\alpha_{2}$ is disjoint from $\beta$), we get $f_{3}\in G$. Since we find that $f_{3}(\alpha_{2})=t_{\alpha_{3}}(\alpha_{2})=t_{\alpha_{2}}^{-1}(\alpha_{3})$, we have $t_{\alpha_{2}}f_{3}(\alpha_{2})=\alpha_{3}$, and therefore we see that
$t_{\alpha_{3}}\in G$
by $t_{\alpha_{2}},t_{\alpha_{2}}f_{3}\in G$. From this, we get $f_{4}=t_{\alpha_{3}}^{-1}f_{3}\in G$ (since $\alpha_{3}$ is disjont from $\beta$). Here, since $\alpha_{3}$ (resp. $\alpha_{4}$) intersects $\alpha_{4}$ (resp. $\beta$) transversely at exactly one point, we find that $f_{4}(\alpha_{3})=t_{\beta}t_{\alpha_{4}}^{-1}(\alpha_{3})=t_{\beta}t_{\alpha_{3}}(\alpha_{4})=t_{\alpha_{3}}t_{\beta}(\alpha_{4})$ (since $\alpha_{3}$ is disjoint from $\beta$). This gives $t_{\alpha_{3}}^{-1}f_{4}(\alpha_{3})=t_{\beta}(\alpha_{4})$, and hence we have
$t_{t_{\beta}(\alpha_{4})}\in G$
by $t_{\alpha_{3}},t_{\alpha_{3}}^{-1}f_{4}\in G$. Furthermore, the following holds:
$G\ni f_{4}$ $=t_{\beta}t_{\alpha_{4}}^{-1}\cdot t_{\alpha_{5}}t_{\alpha_{6}}^{-1}t_{\alpha_{7}}t_{\alpha_{8}}^{-1}\cdots t_{\alpha_{2g-1}}t_{\alpha_{2g}}^{-1}$
$=t_{\beta}t_{\alpha_{4}}^{-1}t_{\beta}^{-1}\cdot t_{\beta}t_{\alpha_{5}}t_{\alpha_{6}}^{-1}t_{\alpha_{7}}t_{\alpha_{8}}^{-1}\cdots t_{\alpha_{2g-1}}t_{\alpha_{2g}}^{-1}$
$=t_{t_{\beta}(\alpha_{4})}f_{5}.$
This means that $f_{5}\in G$. Since $\beta$ and $\alpha_{4}$ are disjont from $\alpha_{6},\ldots,\alpha_{2g}$, we have $f_{5}(t_{\beta}(\alpha_{4}))=t_{\beta}t_{\alpha_{5}}(t_{\beta}(\alpha_{4}))$. Note that $\beta$ is also disjoint from $\alpha_{5}$. Therefore, using the braid relation $t_{\alpha_{5}}t_{\beta}=t_{\beta}t_{\alpha_{5}}$, the relations $t_{\alpha_{5}}(\alpha_{4})=t_{\alpha_{4}}^{-1}(\alpha_{5})$, $t_{\beta}(\alpha_{5})=\alpha_{5}$ and the braid relation $t_{\beta}t_{\alpha_{4}}t_{\beta}=t_{\alpha_{4}}t_{\beta}t_{\alpha_{4}}$, we have
$t_{t_{\beta}(\alpha_{4})}f_{5}(t_{\beta}(\alpha_{4}))$ $=t_{t_{\beta}(\alpha_{4})}t_{\beta}^{2}t_{\alpha_{5}}(\alpha_{4})$
$=t_{t_{\beta}(\alpha_{4})}t_{\beta}^{2}t_{\alpha_{4}}^{-1}(\alpha_{5})$
$=t_{\beta}t_{\alpha_{4}}t_{\beta}^{-1}t_{\beta}^{2}t_{\alpha_{4}}^{-1}(\alpha_{5})$
$=t_{\beta}t_{\alpha_{4}}t_{\beta}t_{\alpha_{4}}^{-1}(\alpha_{5})$
$=t_{\alpha_{4}}t_{\beta}t_{\alpha_{4}}t_{\alpha_{4}}^{-1}(\alpha_{5})$
$=t_{\alpha_{4}}(\alpha_{5}).$
This gives $t_{t_{\alpha_{4}}(\alpha_{5})}=t_{t_{t_{\beta}(\alpha_{4})}f_{5}(t_{\beta}(\alpha_{4}))}(=(t_{t_{\beta}(\alpha_{4})}f_{5})t_{t_{\beta}(\alpha_{4})}(t_{t_{\beta}(\alpha_{4})}f_{5})^{-1})$, and hence we have
$t_{t_{\alpha_{4}}(\alpha_{5})}\in G$
by $t_{t_{\beta}(\alpha_{4})},t_{t_{\beta}(\alpha_{4})}f_{5}\in G$. Since $\alpha_{i}$ is disjoint from $\alpha_{j}$ if $|i-j|\geq 2$, using the relation $t_{\alpha_{6}}^{-1}(\alpha_{5})=t_{\alpha_{5}}(\alpha_{6})$ and the braid relation $t_{\alpha_{5}}t_{\alpha_{4}}t_{\alpha_{5}}=t_{\alpha_{4}}t_{\alpha_{5}}t_{\alpha_{4}}$, we have
$f_{3}(t_{\alpha_{4}}(\alpha_{5}))$ $=t_{\beta}t_{\alpha_{3}}t_{\alpha_{4}}^{-1}t_{\alpha_{5}}t_{\alpha_{6}}^{-1}(t_{\alpha_{4}}(\alpha_{5}))$
$=t_{\beta}t_{\alpha_{3}}t_{\alpha_{4}}^{-1}t_{\alpha_{5}}t_{\alpha_{6}}^{-1}t_{\alpha_{4}}(\alpha_{5})$
$=t_{\beta}t_{\alpha_{3}}t_{\alpha_{4}}^{-1}t_{\alpha_{5}}t_{\alpha_{4}}t_{\alpha_{6}}^{-1}(\alpha_{5})$
$=t_{\beta}t_{\alpha_{3}}t_{\alpha_{4}}^{-1}t_{\alpha_{5}}t_{\alpha_{4}}t_{\alpha_{5}}(\alpha_{6})$
$=t_{\beta}t_{\alpha_{3}}t_{\alpha_{4}}^{-1}t_{\alpha_{4}}t_{\alpha_{5}}t_{\alpha_{4}}(\alpha_{6})$
$=t_{\beta}t_{\alpha_{3}}t_{\alpha_{5}}(\alpha_{6}).$
Since $\beta$ and $\alpha_{3}$ are disjoint from $\alpha_{5}$ and $\alpha_{6}$, we obtain $f_{3}(t_{\alpha_{4}}(\alpha_{5}))=t_{\alpha_{5}}(\alpha_{6})$, and hence we have
$t_{t_{\alpha_{5}}(\alpha_{6})}\in G.$
by $t_{t_{\alpha_{4}}(\alpha_{5})},f_{3}\in G$, and the proof is complete. ∎
Lemma 3.4.
Let $g\geq 4$, and set $F_{2k-1}=t_{\alpha_{3}}t_{\alpha_{5}}\cdots t_{\alpha_{2k-3}}f_{2k-1}$ for $3\leq k\leq g$. We assume that the elements $t_{t_{\alpha_{2i-1}}(\alpha_{2i})},F_{2i-1},F_{2i+1}$ are contained in $G$ for some $i$ such that $13\leq i\leq g-$ 1, then we have
  1. (1)
    $t_{t_{\alpha_{2i}}(\alpha_{2i+1})},t_{t_{\alpha_{2i+1}}(\alpha_{2i+2})}\in G$ if $13\leq i\leq g-$ 1,
  2. (2)
    $F_{2i+3}\in G$ if $i\leq g-2$, and
  3. (3)
    $t_{\alpha_{2g}}\in G$ if $i=g-1$.
Proof.
First, we show the former part of Lemma 3.4 (1) (i.e., the element $t_{t_{\alpha_{2i}}(\alpha_{2i+1})}$ is in $G$ for $13\leq i\leq g-$ 1). Note that $\alpha_{2i-1}$ and $\alpha_{2i}$ are disjoint from $\alpha_{j}$ for $j\leq 2i-3$ or $2i+2\leq j$. Moreover, if $3\leq i$ then $\alpha_{2i-1},\alpha_{2i},\alpha_{2i+1}$ are disjoint from $\beta$. From these, we have
$f_{2i+1}(t_{\alpha_{2i-1}}(\alpha_{2i}))$ $=t_{\beta}t_{\alpha_{2i+1}}t_{\alpha_{2i+2}}^{-1}\cdots t_{\alpha_{2g-1}}t_{\alpha_{2g}}^{-1}(t_{\alpha_{2i-1}}(\alpha_{2i}))$
$=t_{\beta}t_{\alpha_{2i+1}}(t_{\alpha_{2i-1}}(\alpha_{2i}))$
$=t_{\alpha_{2i+1}}(t_{\alpha_{2i-1}}(\alpha_{2i})).$
Since $\alpha_{2i-1},\alpha_{2i},\alpha_{2i+1}$ are disjoint from $\alpha_{3},\alpha_{5},\ldots,\alpha_{2i-3}$, we obtain
$F_{2i+1}(t_{\alpha_{2i-1}}(\alpha_{2i}))$ $=t_{\alpha_{3}}t_{\alpha_{5}}\cdots t_{\alpha_{2i-3}}t_{\alpha_{2i-1}}f_{2i+1}(t_{\alpha_{2i-1}}(\alpha_{2i}))$
$=t_{\alpha_{3}}t_{\alpha_{5}}\cdots t_{\alpha_{2i-3}}t_{\alpha_{2i-1}}(t_{\alpha_{2i+1}}(t_{\alpha_{2i-1}}(\alpha_{2i})))$
$=t_{\alpha_{2i-1}}(t_{\alpha_{2i+1}}(t_{\alpha_{2i-1}}(\alpha_{2i})))$
$=t_{\alpha_{2i-1}}t_{\alpha_{2i+1}}t_{\alpha_{2i-1}}(\alpha_{2i}).$
From the above arguments, the braid relation $t_{\alpha_{2i-1}}t_{\alpha_{2i}}t_{\alpha_{2i-1}}=t_{\alpha_{2i}}t_{\alpha_{2i-1}}t_{\alpha_{2i}}$ and the relation $t_{\alpha_{2i}}t_{\alpha_{2i+1}}(\alpha_{2i})=\alpha_{2i+1}$, we have
$t_{t_{\alpha_{2i-1}}(\alpha_{2i})}F_{2i+1}(t_{\alpha_{2i-1}}(\alpha_{2i}))$ $=t_{t_{\alpha_{2i-1}}(\alpha_{2i})}(t_{\alpha_{2i-1}}t_{\alpha_{2i+1}}t_{\alpha_{2i-1}}(\alpha_{2i}))$
$=t_{\alpha_{2i-1}}t_{\alpha_{2i}}t_{\alpha_{2i-1}}^{-1}t_{\alpha_{2i-1}}t_{\alpha_{2i+1}}t_{\alpha_{2i-1}}(\alpha_{2i})$
$=t_{\alpha_{2i-1}}t_{\alpha_{2i}}t_{\alpha_{2i+1}}t_{\alpha_{2i-1}}(\alpha_{2i})$
$=t_{\alpha_{2i-1}}t_{\alpha_{2i}}t_{\alpha_{2i-1}}t_{\alpha_{2i+1}}(\alpha_{2i})$
$=t_{\alpha_{2i}}t_{\alpha_{2i-1}}t_{\alpha_{2i}}t_{\alpha_{2i+1}}(\alpha_{2i})$
$=t_{\alpha_{2i}}t_{\alpha_{2i-1}}(\alpha_{2i+1})$
$=t_{\alpha_{2i}}(\alpha_{2i+1}).$
This gives $t_{t_{\alpha_{2i}}(\alpha_{2i+1})}=t_{t_{t_{\alpha_{2i-1}}(\alpha_{2i})}F_{2i+1}(t_{\alpha_{2i-1}}(\alpha_{2i}))}$, and hence the element $t_{t_{\alpha_{2i}}(\alpha_{2i+1})}$ is in $G$ by $t_{t_{\alpha_{2i-1}}(\alpha_{2i})},t_{t_{\alpha_{2i-1}}(\alpha_{2i})}F_{2i+1}\in G$ (since $t_{t_{\alpha_{2i-1}}(\alpha_{2i})}$ and $F_{2i+1}$ are in $G$).
Second, we show the latter part of Lemma 3.4 (1) (i.e., the element $t_{t_{\alpha_{2i+1}}(\alpha_{2i+2})}$ is in $G$ for $13\leq i\leq g-$ 1). Note that $\alpha_{2i}$ and $\alpha_{2i+1}$ are disjoint from $\alpha_{j}$ for $2i+3\leq j$. Moreover, if $3\leq i$ then $\alpha_{2i-1},\alpha_{2i},\alpha_{2i+1},\alpha_{2i+2}$ are disjoint from $\beta$. From these and the relation $t_{\alpha_{2i+2}}^{-1}(\alpha_{2i+1})=t_{\alpha_{2i+1}}(\alpha_{2i+2})$, we have
$f_{2i-1}(t_{\alpha_{2i}}(\alpha_{2i+1}))$ $=t_{\beta}t_{\alpha_{2i-1}}t_{\alpha_{2i}}^{-1}\cdots t_{\alpha_{2g-1}}t_{\alpha_{2g}}^{-1}(t_{\alpha_{2i}}(\alpha_{2i+1}))$
$=t_{\beta}t_{\alpha_{2i-1}}t_{\alpha_{2i}}^{-1}t_{\alpha_{2i+1}}t_{\alpha_{2i+2}}^{-1}(t_{\alpha_{2i}}(\alpha_{2i+1}))$
$=t_{\alpha_{2i-1}}t_{\alpha_{2i}}^{-1}t_{\alpha_{2i+1}}t_{\alpha_{2i+2}}^{-1}t_{\alpha_{2i}}(\alpha_{2i+1})$
$=t_{\alpha_{2i-1}}t_{\alpha_{2i}}^{-1}t_{\alpha_{2i+1}}t_{\alpha_{2i}}t_{\alpha_{2i+2}}^{-1}(\alpha_{2i+1})$
$=t_{\alpha_{2i-1}}t_{\alpha_{2i}}^{-1}t_{\alpha_{2i+1}}t_{\alpha_{2i}}t_{\alpha_{2i+1}}(\alpha_{2i+2}).$
Moreover, using the braid relation $t_{\alpha_{2i+1}}t_{\alpha_{2i}}t_{\alpha_{2i+1}}=t_{\alpha_{2i}}t_{\alpha_{2i+1}}t_{\alpha_{2i}}$, the following holds:
$t_{\alpha_{2i-1}}t_{\alpha_{2i}}^{-1}t_{\alpha_{2i+1}}t_{\alpha_{2i}}t_{\alpha_{2i+1}}$ $=t_{\alpha_{2i-1}}t_{\alpha_{2i}}^{-1}t_{\alpha_{2i}}t_{\alpha_{2i+1}}t_{\alpha_{2i}}$
$=t_{\alpha_{2i-1}}t_{\alpha_{2i+1}}t_{\alpha_{2i}}$
$=t_{\alpha_{2i+1}}t_{\alpha_{2i-1}}t_{\alpha_{2i}}.$
Since $\alpha_{2i-1}$ and $\alpha_{2i}$ are disjoint from $\alpha_{2i+2}$, the above arguments give
$f_{2i-1}(t_{\alpha_{2i}}(\alpha_{2i+1}))=t_{\alpha_{2i+1}}t_{\alpha_{2i-1}}t_{\alpha_{2i}}(\alpha_{2i+2})=t_{\alpha_{2i+1}}(\alpha_{2i+2}),$
and moreover, since $\alpha_{2i+1}$ and $\alpha_{2i+2}$ are disjoint from $\alpha_{3},\alpha_{5},\ldots,\alpha_{2i-3}$, we have
$F_{2i-1}(t_{\alpha_{2i}}(\alpha_{2i+1}))$ $=t_{\alpha_{3}}t_{\alpha_{5}}\cdots t_{\alpha_{2i-3}}f_{2i-1}(t_{\alpha_{2i}}(\alpha_{2i+1}))$
$=t_{\alpha_{3}}t_{\alpha_{5}}\cdots t_{\alpha_{2i-3}}(t_{\alpha_{2i+1}}(\alpha_{2i+2}))$
$=t_{\alpha_{2i+1}}(\alpha_{2i+2}).$
This gives $t_{t_{\alpha_{2i+1}}(\alpha_{2i+2})}=t_{F_{2i-1}(t_{\alpha_{2i}}(\alpha_{2i+1}))}$, and hence the element $t_{t_{\alpha_{2i+1}}(\alpha_{2i+2})}$ is in $G$ by $F_{2i-1},t_{t_{\alpha_{2i}}(\alpha_{2i+1})}\in G$.
Third, we show Lemma 3.4 (2) (i.e., the element $F_{2i+3}$ is in $G$ for $i\leq g-2$). Since $\alpha_{2i+1}$ and $\alpha_{2i+2}$ are disjoint from $\beta$ if $3\leq i$, by the relation $t_{\alpha_{2i+1}}t_{\alpha_{2i+2}}^{-1}t_{\alpha_{2i+1}}^{-1}=t_{t_{\alpha_{2i+1}}(\alpha_{2i+2})}^{-1}$ we obtain
$f_{2i+1}$ $=t_{\beta}t_{\alpha_{2i+1}}t_{\alpha_{2i+2}}^{-1}t_{\alpha_{2i+3}}t_{\alpha_{2i+4}}^{-1}\cdots t_{\alpha_{2g-1}}t_{\alpha_{2g}}^{-1}$
$=t_{\alpha_{2i+1}}t_{\alpha_{2i+2}}^{-1}t_{\beta}t_{\alpha_{2i+3}}t_{\alpha_{2i+4}}^{-1}\cdots t_{\alpha_{2g-1}}t_{\alpha_{2g}}^{-1}$
$=t_{\alpha_{2i+1}}t_{\alpha_{2i+2}}^{-1}t_{\alpha_{2i+1}}^{-1}t_{\alpha_{2i+1}}\cdot t_{\beta}t_{\alpha_{2i+3}}t_{\alpha_{2i+4}}^{-1}\cdots t_{\alpha_{2g-1}}t_{\alpha_{2g}}^{-1}$
$=t_{t_{\alpha_{2i+1}}(\alpha_{2i+2})}^{-1}t_{\alpha_{2i+1}}\cdot f_{2i+3}.$
Since $\alpha_{2i+1}$ and $\alpha_{2i+2}$ are disjoint from $\alpha_{j}$ for $j\leq 2i-1$, we have
$G\ni F_{2i+1}$ $=t_{\alpha_{3}}t_{\alpha_{5}}\cdots t_{\alpha_{2i-3}}t_{\alpha_{2i-1}}f_{2i+1}$
$=t_{\alpha_{3}}t_{\alpha_{5}}\cdots t_{\alpha_{2i-3}}t_{\alpha_{2i-1}}\cdot t_{t_{\alpha_{2i+1}}(\alpha_{2i+2})}^{-1}t_{\alpha_{2i+1}}\cdot f_{2i+3}$
$=t_{t_{\alpha_{2i+1}}(\alpha_{2i+2})}^{-1}\cdot t_{\alpha_{3}}t_{\alpha_{5}}\cdots t_{\alpha_{2i-3}}t_{\alpha_{2i-1}}\cdot t_{\alpha_{2i+1}}\cdot f_{2i+3}$
$=t_{t_{\alpha_{2i+1}}(\alpha_{2i+2})}^{-1}F_{2i+3}.$
By $t_{t_{\alpha_{2i+1}}(\alpha_{2i+2})}\in G$ (i.e., Lemma 3.4 (1)), the element $F_{2i+3}$ is in $G$.
Finally, we show Lemma 3.4 (3) (i.e., the element $t_{\alpha_{2g}}$ is in $G$ for $i=g-1$). If $i=g-1$, then we have $F_{2i+1}=F_{2g-1}=t_{\alpha_{3}}t_{\alpha_{5}}\cdots t_{\alpha_{2g-3}}f_{2g-1}\in G$. Since $\alpha_{2g-1}$ and $\alpha_{2g}$ are disjoint from $\beta,\alpha_{3},\alpha_{5},\ldots,\alpha_{2g-3}$, we see that the following holds:
$F_{2g-1}=t_{\alpha_{3}}t_{\alpha_{5}}\cdots t_{\alpha_{2g-3}}f_{2g-1}$ $=t_{\alpha_{3}}t_{\alpha_{5}}\cdots t_{\alpha_{2g-3}}t_{\beta}t_{\alpha_{2g-1}}t_{\alpha_{2g}}^{-1}$
$=t_{\alpha_{3}}t_{\alpha_{5}}\cdots t_{\alpha_{2g-3}}t_{\beta}\cdot t_{\alpha_{2g-1}}t_{\alpha_{2g}}^{-1}t_{\alpha_{2g-1}}^{-1}t_{\alpha_{2g-1}}$
$=t_{\alpha_{2g-1}}t_{\alpha_{2g}}^{-1}t_{\alpha_{2g-1}}^{-1}\cdot t_{\alpha_{3}}t_{\alpha_{5}}\cdots t_{\alpha_{2g-3}}t_{\beta}t_{\alpha_{2g-1}}$
$=t_{t_{\alpha_{2g-1}(\alpha_{2g})}}\cdot t_{\alpha_{3}}t_{\alpha_{5}}\cdots t_{\alpha_{2g-3}}t_{\beta}t_{\alpha_{2g-1}}.$
By $F_{2g-1}=t_{t_{\alpha_{2g-1}(\alpha_{2g})}}\cdot t_{\alpha_{3}}t_{\alpha_{5}}\cdots t_{\alpha_{2g-3}}t_{\beta}t_{\alpha_{2g-1}}\in G$ and $t_{t_{\alpha_{2g-1}(\alpha_{2g})}}\in G$ (i.e., Lemma 3.4 (1)), we have $t_{\alpha_{3}}t_{\alpha_{5}}\cdots t_{\alpha_{2g-3}}t_{\beta}t_{\alpha_{2g-1}}\in G$. Therefore, the element $t_{\alpha_{2g}}$ is in $G$ by $G\ni F_{2g-1}=(t_{\alpha_{3}}t_{\alpha_{5}}\cdots t_{\alpha_{2g-3}}t_{\beta}t_{\alpha_{2g-1}})t_{\alpha_{2g}}^{-1}$, and the lemma follows. ∎
Proof of Theorem 3.2.
We note that we have seen that the elements $t_{\alpha_{1}}$, $t_{\alpha_{2}}$, $t_{\alpha_{3}}$, $t_{t_{\alpha_{4}}(\alpha_{5})}$, $t_{t_{\alpha_{5}}(\alpha_{6})}$ and $f_{5}$ are contained in $G$ by Lemma 3.3.
Suppose that $g=3$. Since $\alpha_{5}$ and $\alpha_{6}$ are disjoint from $\beta$, we have
$G\ni f_{5}=t_{\beta}t_{\alpha_{5}}t_{\alpha_{6}}^{-1}=t_{\alpha_{5}}t_{\alpha_{6}}^{-1}t_{\alpha_{5}}^{-1}t_{\beta}t_{\alpha_{5}}=t_{t_{\alpha_{5}}(\alpha_{6})}^{-1}\cdot t_{\beta}t_{\alpha_{5}},$
and hence the element $t_{\beta}t_{\alpha_{5}}$ is in $G$ by $t_{t_{\alpha_{5}}(\alpha_{6})}\in G$. Moreover, by $G\ni f_{5}=(t_{\beta}t_{\alpha_{5}})t_{\alpha_{6}}^{-1}$, we have $t_{\alpha_{6}}\in G$. Thus, since $t_{\alpha_{5}}(\alpha_{6})=t_{\alpha_{6}}^{-1}(\alpha_{5})$, we see that $t_{t_{\alpha_{5}}(\alpha_{6})}=t_{t_{\alpha_{6}^{-1}}(\alpha_{5})}(=t_{\alpha_{6}}^{-1}t_{\alpha_{5}}t_{\alpha_{6}}$), and hence the element $t_{\alpha_{5}}$ is in $G$ by $t_{t_{\alpha_{5}}(\alpha_{6})},t_{\alpha_{6}}\in G$. By applying this argument, with $t_{\alpha_{6}},t_{t_{\alpha_{5}}(\alpha_{6})}\in G$ replaced by $t_{\alpha_{5}},t_{t_{\alpha_{4}}(\alpha_{5})}\in G$, we find that the element $t_{\alpha_{4}}$ is in $G$. Therefore, the elements $t_{\alpha_{1}},t_{\alpha_{2}},\ldots,t_{\alpha_{6}}$ are contained in $G$, and hence we see that $t_{\beta}$ is in $G$ by $f_{1}=t_{\beta}t_{\alpha_{1}}t_{\alpha_{2}}^{-1}\cdots t_{\alpha_{5}}t_{\alpha_{6}}^{-1}\in G$. Since $\mathcal{M}_{g}$ is generated by $t_{\alpha_{1}},t_{\alpha_{2}},\ldots,t_{\alpha_{6}},t_{\beta}$ if $g=3$, we see that $G=\mathcal{M}_{g}$.
Suppose that $g\geq 4$. Since $\alpha_{5}$ and $\alpha_{6}$ are disjoint from $\beta$, we have
$G\ni f_{5}$ $=t_{\beta}t_{\alpha_{5}}t_{\alpha_{6}}^{-1}t_{\alpha_{7}}t_{\alpha_{8}}^{-1}\cdots t_{\alpha_{2g-1}}t_{\alpha_{2g}}^{-1}$
$=t_{\alpha_{5}}t_{\alpha_{6}}^{-1}t_{\alpha_{5}}^{-1}t_{\alpha_{5}}\cdot t_{\beta}t_{\alpha_{7}}t_{\alpha_{8}}^{-1}\cdots t_{\alpha_{2g-1}}t_{\alpha_{2g}}^{-1}$
$=t_{t_{\alpha_{5}}(\alpha_{6})}^{-1}t_{\alpha_{5}}f_{7}.$
By $t_{t_{\alpha_{5}}(\alpha_{6})}\in G$, the element $t_{\alpha_{5}}f_{7}$ is contained in $G$. Moreover, since $t_{\alpha_{3}}$ and $f_{5}$ are in $G$, we have $F_{5}=t_{\alpha_{3}}f_{5}\in G$ and $F_{7}=t_{\alpha_{3}}t_{\alpha_{5}}f_{7}\in G$. Therefore, the condition of Lemma 3.4 is satisfied for the case where $i=3$. By applying Lemma 3.4 repeatedly, we have $t_{t_{\alpha_{j}}(\alpha_{j+1})}\in G$ for $j=6,7,\ldots,2g-1$ and $t_{\alpha_{2g}}\in G$. Here, by $t_{\alpha_{2g-1}(\alpha_{2g})}=t_{\alpha_{2g}}^{-1}(\alpha_{2g-1})$, we have $t_{t_{\alpha_{2g-1}}(\alpha_{2g})}=t_{t_{\alpha_{2g}}^{-1}(\alpha_{2g-1})}(=t_{\alpha_{2g}}^{-1}t_{\alpha_{2g-1}}t_{\alpha_{2g}})$, and hence the element $t_{\alpha_{2g-1}}$ is in $G$ by $t_{t_{\alpha_{2g-1}}(\alpha_{2g})},t_{\alpha_{2g}}\in G$. By repeating this argument, we find that the elements $t_{\alpha_{4}},t_{\alpha_{5}},\ldots,t_{\alpha_{2g}}$ are contained in $G$. Moreover, from the definition of $f_{1}$, we have $t_{\beta}\in G$, and hence $G=\mathcal{M}_{g}$.
This finishes the proof. ∎
Theorem 1.4 follows from Theorem 3.5 below.
Theorem 3.5.
Let $H^{\prime}$ be the subgroup of $\mathcal{H}_{g}$ generated by $\mathcal{R}$, which is defined in Lemma 2.7, and $r_{\pi}\mathcal{R}r_{\pi}^{-1}$, where $r_{\pi}$ is the rotation defined in Section 2. Then, $H^{\prime}=\mathcal{H}_{g}$ for $g\geq 2$, and $\mathcal{H}_{1}$ is generated by $t_{\alpha_{1}}$ and $t_{\alpha_{2}}$.
Proof.
Suppose that $g\geq 2$. From Lemma 2.7, we have $t_{\alpha_{2g+1}}^{-1}=\mathcal{R}^{2g-1}\in H^{\prime}$, and therefore $(t_{\alpha_{1}}^{2}t_{\alpha_{2}}t_{\alpha_{3}}\cdots t_{\alpha_{2g-1}})^{g-1}=\mathcal{R}t_{\alpha_{2g+1}}\in H^{\prime}$. From these and the chain relation $(t_{\alpha_{1}}^{2}t_{\alpha_{2}}t_{\alpha_{3}}\cdots t_{\alpha_{2g-1}})^{2g-1}=t_{\alpha_{2g+1}}^{2}$ (see the proof of Lemma 2.7), we obtain the following:
$t_{\alpha_{1}}^{2}t_{\alpha_{2}}t_{\alpha_{3}}\cdots t_{\alpha_{2g-1}}=t_{\alpha_{2g+1}}^{2}(t_{\alpha_{1}}^{2}t_{\alpha_{2}}t_{\alpha_{3}}\cdots t_{\alpha_{2g-1}})^{-2(g-1)}\in H^{\prime}.$
Therefore, by $H^{\prime}\ni(r_{\pi}\mathcal{R}r_{\pi}^{-1})^{2g-1}=r_{\pi}\mathcal{R}^{2g-1}r_{\pi}^{-1}=r_{\pi}t_{\alpha_{2g+1}}^{-1}r_{\pi}^{-1}=t_{r_{\pi}(\alpha_{2g+1})}^{-1}=t_{\alpha_{1}}^{-1}$, we have $t_{\alpha_{1}},t_{\alpha_{1}}t_{\alpha_{2}}t_{\alpha_{3}}\cdots t_{\alpha_{2g-1}}\in H^{\prime}$. It is easy to check that $t_{\alpha_{1}}t_{\alpha_{2}}t_{\alpha_{3}}\cdots t_{\alpha_{2g-1}}(\alpha_{j})=\alpha_{j+1}$ for $j=1,2,\ldots,2g-2$, and we have $t_{\alpha_{j+1}}=t_{f(\alpha_{j})}=ft_{\alpha_{j}}f^{-1}$, where $f=t_{\alpha_{1}}t_{\alpha_{2}}t_{\alpha_{3}}\cdots t_{\alpha_{2g-1}}$. From this and $t_{\alpha_{1}}\in H^{\prime}$, we obtain that the elements $t_{\alpha_{2}},t_{\alpha_{3}},\ldots,t_{\alpha_{2g-1}}$ are contained in $H^{\prime}$. Here, we see that
$r_{\pi}\mathcal{R}r_{\pi}^{-1}$ $=r_{\pi}(t_{\alpha_{1}}^{2}t_{\alpha_{2}}t_{\alpha_{3}}\cdots t_{\alpha_{2g-1}})^{g-1}t_{\alpha_{2g+1}}^{-1}r_{\pi}^{-1}$
$=((r_{\pi}t_{\alpha_{1}}r_{\pi}^{-1})^{2}\cdot r_{\pi}t_{\alpha_{2}}r_{\pi}^{-1}\cdot r_{\pi}t_{\alpha_{3}}r_{\pi}^{-1}\cdots r_{\pi}t_{\alpha_{2g-1}}r_{\pi}^{-1})^{g-1}r_{\pi}t_{\alpha_{2g+1}}^{-1}r_{\pi}^{-1}$
$=(t_{r_{\pi}(\alpha_{1})}^{2}t_{r_{\pi}(\alpha_{2})}t_{r_{\pi}(\alpha_{3})}\cdots t_{r_{\pi}(\alpha_{2g-1})})^{g-1}t_{r_{\pi}(\alpha_{2g+1})}^{-1}$
$=(t_{\alpha_{2g+1}}^{2}t_{\alpha_{2g}}t_{\alpha_{2g-1}}\cdots t_{\alpha_{3}})^{g-1}t_{\alpha_{1}}^{-1},$
and hence we have $(t_{\alpha_{2g+1}}^{2}t_{\alpha_{2g}}t_{\alpha_{2g-1}}\cdots t_{\alpha_{3}})^{g-1}=r_{\pi}\mathcal{R}r_{\pi}^{-1}t_{\alpha_{1}}\in H^{\prime}$ by $t_{\alpha_{1}}\in H^{\prime}$. We set $h=t_{\alpha_{2g+1}}^{2}t_{\alpha_{2g}}t_{\alpha_{2g-1}}\cdots t_{\alpha_{3}}$, and so $h^{g-1}$ is in $H^{\prime}$. It is easily seen that $h^{-1}(\alpha_{i})=\alpha_{i+1}$ for $i=3,4,\ldots,2g-1$, and hence $h^{-(g-1)}(\alpha_{g+1})=\alpha_{2g}$. Therefore, we have $t_{\alpha_{2g}}\in H^{\prime}$ by $t_{\alpha_{g+1}},h^{g-1}\in H^{\prime}$ and the relation $t_{h^{-(g-1)}(\alpha_{g+1})}=h^{-(g-1)}t_{\alpha_{g+1}}h^{g-1}$. Summarizing, the elements $t_{\alpha_{1}},t_{\alpha_{2}},\ldots,t_{\alpha_{2g+1}}$ are contained in $H^{\prime}$, and hence $H^{\prime}=\mathcal{H}_{g}$ for $g\geq 2$. By Theorem 2.1, $\mathcal{H}_{1}$ is generated by $t_{\alpha_{1}}$ and $t_{\alpha_{2}}$. This completes the proof. ∎
4. Proofs of Theorems 1.2 and 1.3
In this section, Theorems 1.2 and 1.3 in the introduction. Theorem 1.2 follows from the Thorems below.
Theorem 4.1.
Let $g\geq 9$, and let $G_{1}$ be the subgroup of $\mathcal{M}_{g}$ generated by $\rho_{n}$ and $(t_{b_{3}}t_{a_{5}}^{-1}t_{c_{6}})\rho_{n}(t_{b_{3}}t_{a_{5}}^{-1}t_{c_{6}})^{-1}$, where $\rho_{n}$ is defined in Lemma 2.5. Then, $G_{1}=\mathcal{M}_{g}$.
Theorem 4.2.
For $g\geq 3$, let $G_{2}$ be the subgroup of $\mathcal{M}_{g}$ generated by $\rho_{n}$, $t_{a_{2}}\rho_{n}t_{a_{2}}^{-1}$ and $(t_{a_{3}}t_{b_{3}}t_{c_{2}})\rho_{n}(t_{a_{3}}t_{b_{3}}t_{c_{2}})^{-1}$, where $\rho_{n}$ is defined in Lemma 2.5. Then, $G_{2}=\mathcal{M}_{g}$.
Theorem 1.3 follows from the Thorems below.
Theorem 4.3.
Let $g\geq 8$, and let $G_{3}$ be the subgroup of $\mathcal{M}_{g}$ generated by $\rho$ and $(t_{b_{2}}t_{a_{4}}^{-1}t_{c_{5}})\rho(t_{b_{2}}t_{a_{4}}^{-1}t_{c_{5}})^{-1}$, where $\rho$ is defined in Lemma 2.6. Then, $G_{3}=\mathcal{M}_{g}$.
Theorem 4.4.
For $g\geq 3$, let $G_{4}$ be the subgroup of $\mathcal{M}_{g}$ generated by $\rho^{\prime}$, $t_{a_{2}}\rho^{\prime}t_{a_{2}}^{-1}$ and $(t_{a_{3}}t_{b_{3}}t_{c_{2}})\rho^{\prime}(t_{a_{3}}t_{b_{3}}t_{c_{2}})^{-1}$, where $\rho^{\prime}$ is defined in Lemma 2.6. Then, $G_{4}=\mathcal{M}_{g}$.
Lemmas 4.54.7 below are used to show Theorems 4.14.4. Recall that if a Dehn twists $t_{c}$ about a simple closed curve $c$ and an element $f$ are in a subgroup $G$ of $\mathcal{M}_{g}$, then $t_{f}(c)$ is also in $G$ by the relation $t_{f(c)}=ft_{c}f^{-1}$. Moreover, if the product $t_{c}t_{d}^{-1}$ of two Dehn twists $t_{c}$ and $t_{d}$ about two simple closed curve $c$ and $d$ and an element $f$ are in $G$, then $t_{f(c)}t_{f(d)}^{-1}$ is also in $G$ by $t_{f(c)}t_{f(d)}^{-1}=ft_{c}f^{-1}ft_{d}^{-1}f^{-1}=ft_{c}t_{d}^{-1}f^{-1}$. We will repeatedly use these facts in the proofs of Lemmas 4.54.7 and Theorems 4.14.4.
Lemma 4.5.
Let $g\geq 3$, and let $G^{\prime}$ be a subgroup of $\mathcal{M}_{g}$ which contains
$t_{a_{k}}t_{a_{k+1}}^{-1},\ t_{a_{k}}t_{a_{k+2}}^{-1},\ t_{a_{k}}t_{b_{k+1}}^{-1},\ t_{a_{k}}t_{c_{k}}^{-1},\ t_{a_{k+1}}t_{c_{k+1}}^{-1},\ t_{a_{k}}t_{c_{k+1}}^{-1},\ t_{a_{k}}t_{b_{k+2}}^{-1}$
for some integer $k$ such that $21\leq k\leq g-$ 2. Then, the element $t_{a_{k+2}}$ is in $G^{\prime}$.
Proof.
Let $d_{1}$ and $d_{2}$ be the simple closed curves in $\Sigma_{g}$ as in Figure 5. Then, it is easy to check that
$t_{b_{k+1}}t_{a_{k}}^{-1}\cdot t_{c_{k}}t_{a_{k}}^{-1}\cdot t_{a_{k}}t_{a_{k+1}}^{-1}\cdot t_{c_{k+1}}t_{a_{k}}^{-1}(b_{k+1},a_{k})$ $=(d_{1},a_{k})$
$t_{b_{k+2}}t_{a_{k}}^{-1}\cdot t_{c_{k+1}}t_{a_{k}}^{-1}\cdot t_{a_{k+2}}t_{a_{k}}^{-1}\cdot t_{b_{k+2}}t_{a_{k}}^{-1}(d_{1},a_{k})$ $=(d_{2},a_{k}).$
When we set $\phi_{1}=t_{b_{k+1}}t_{a_{k}}^{-1}\cdot t_{c_{k}}t_{a_{k}}^{-1}\cdot t_{a_{k}}t_{a_{k+1}}^{-1}\cdot t_{c_{k+1}}t_{a_{k}}^{-1}$ and $\phi_{2}=t_{b_{k+2}}t_{a_{k}}^{-1}\cdot t_{c_{k+1}}t_{a_{k}}^{-1}\cdot t_{a_{k+2}}t_{a_{k}}^{-1}\cdot t_{b_{k+2}}t_{a_{k}}^{-1}$, by $\phi_{1},\phi_{2},t_{b_{k+1}}t_{a_{k}}^{-1}\in G^{\prime}$ we see that
$t_{d_{1}}t_{a_{k}}^{-1}\in G^{\prime},$
$t_{d_{2}}t_{a_{k}}^{-1}=\phi_{2}\cdot t_{d_{1}}t_{a_{k}}^{-1}\cdot\phi_{2}^{-1}\in G^{\prime}.$
Moreover, by $t_{a_{k}}t_{c_{k}}^{-1}\in G^{\prime}$, we have $t_{d_{2}}t_{c_{k}}^{-1}=t_{d_{2}}t_{a_{k}}^{-1}\cdot t_{a_{k}}t_{c_{k}}^{-1}\in G^{\prime}$. Therefore, by the lantern relation $t_{a_{k}}t_{c_{k}}t_{c_{k+1}}t_{a_{k+2}}=t_{a_{k+1}}t_{d_{1}}t_{d_{2}}$, we have
$t_{a_{k+2}}=t_{a_{k+1}}t_{c_{k+1}}^{-1}\cdot t_{d_{1}}t_{a_{k}}^{-1}\cdot t_{d_{2}}t_{c_{k}}^{-1}\in G^{\prime}.$
Figure 5. The simple closed curve $d_{1}$ and $d_{2}$ on $\Sigma_{g}$.
Lemma 4.6.
Let $g\geq k+2$, where $k$ is a positive integer, and let $f$ be an element in $\mathcal{M}_{g}$ satisfying $f(a_{k},b_{k},c_{k})=(a_{k+1},b_{k+1},c_{k+1})$ and $f(a_{k+1},b_{k+1})=(a_{k+2},b_{k+2})$. We denote by $G^{\prime}$ the subgroup of $\mathcal{M}_{g}$ generated by $f$, $t_{a_{k}}t_{a_{k+1}}^{-1}$, $t_{b_{k}}t_{b_{k+1}}^{-1}$ and $t_{c_{k}}t_{c_{k+1}}^{-1}$. Then, the elements $t_{a_{k}},t_{b_{k}},t_{c_{k}}$ are in $G^{\prime}$.
Proof.
By $f,t_{a_{k}}t_{a_{k+1}}^{-1},t_{b_{k}}t_{b_{k+1}}^{-1}\in G^{\prime}$, the following holds:
$t_{a_{k+1}}t_{a_{k+2}}^{-1}\in G^{\prime},$
$t_{b_{k+1}}t_{b_{k+2}}^{-1}\in G^{\prime}.$
Since it is easy to check that $t_{a_{k}}t_{a_{k+1}}^{-1}\cdot t_{b_{k}}t_{b_{k+1}}^{-1}(a_{k+1},a_{k+2})=(b_{k+1},a_{k+2})$, by $t_{a_{k}}t_{a_{k+1}}^{-1},t_{b_{k}}t_{b_{k+1}}^{-1},t_{a_{k+1}}t_{a_{k+2}}^{-1}\in G^{\prime}$, we have
$t_{b_{k+1}}t_{a_{k+2}}^{-1}\in G^{\prime},$
$t_{b_{k}}t_{a_{k+2}}^{-1}=t_{b_{k}}t_{b_{k+1}}^{-1}\cdot t_{b_{k+1}}t_{a_{k+2}}^{-1}\in G^{\prime}.$
Similarly, since it is easily seen that $t_{b_{k}}t_{a_{k+2}}^{-1}\cdot t_{c_{k}}t_{c_{k+1}}^{-1}(b_{k},a_{k+2})=(c_{k},a_{k+2})$, by $t_{b_{k}}t_{a_{k+2}}^{-1},t_{c_{k}}t_{c_{k+1}}^{-1}\in G^{\prime}$, we have
$t_{c_{k}}t_{a_{k+2}}^{-1}\in G^{\prime}.$
We pick up elements in the above arguments to prove Lemma 4.6:
$t_{a_{k}}t_{a_{k+1}}^{-1},\ t_{a_{k+1}}t_{a_{k+2}}^{-1},\ t_{b_{k+1}}t_{a_{k+2}}^{-1},\ t_{c_{k}}t_{a_{k+2}}^{-1}\in G^{\prime}.$
Using these elements, we have
$t_{a_{k}}t_{a_{k+2}}^{-1}$ $=t_{a_{k}}t_{a_{k+1}}^{-1}\cdot t_{a_{k+1}}t_{a_{k+2}}^{-1}\in G^{\prime},$
$t_{a_{k}}t_{b_{k+1}}^{-1}$ $=t_{a_{k}}t_{a_{k+1}}^{-1}\cdot t_{a_{k+1}}t_{a_{k+2}}^{-1}\cdot(t_{b_{k+1}}t_{a_{k+2}}^{-1})^{-1}\in G^{\prime},$
$t_{a_{k}}t_{c_{k}}^{-1}$ $=t_{a_{k}}t_{a_{k+1}}^{-1}\cdot t_{a_{k+1}}t_{a_{k+2}}^{-1}\cdot(t_{c_{k}}t_{a_{k+2}}^{-1})^{-1}\in G^{\prime},$
$t_{a_{k+1}}t_{c_{k+1}}^{-1}$ $=t_{f(a_{k})}t_{f(c_{k})}^{-1}=ft_{a_{k}}t_{c_{k}}^{-1}f^{-1}\in G^{\prime},$
$t_{a_{k}}t_{c_{k+1}}^{-1}$ $=t_{a_{k}}t_{a_{k+1}}^{-1}\cdot t_{a_{k+1}}t_{c_{k+1}}^{-1}\in G^{\prime}.$
Moreover, since we have $t_{a_{k+1}}t_{b_{k+2}}^{-1}\in G^{\prime}$ by the assumption that $f,t_{a_{k}}t_{b_{k+1}}^{-1}\in G^{\prime}$, the element $t_{a_{k}}t_{b_{k+2}}^{-1}$ is in $G^{\prime}$ by $t_{a_{k}}t_{b_{k+2}}^{-1}=t_{a_{k}}t_{a_{k+1}}^{-1}\cdot t_{a_{k+1}}t_{b_{k+2}}^{-1}$. Summarizing, we have
$t_{a_{k}}t_{a_{k+1}}^{-1},\ t_{a_{k}}t_{a_{k+2}}^{-1},\ t_{a_{k}}t_{b_{k+1}}^{-1},\ t_{a_{k}}t_{c_{k}}^{-1},\ t_{a_{k+1}}t_{c_{k+1}}^{-1},\ t_{a_{k}}t_{c_{k+1}}^{-1},\ t_{a_{k}}t_{b_{k+2}}^{-1}\in G^{\prime},$
and we see that $t_{a_{k+2}}\in G^{\prime}$ from Lemma 4.5. By $t_{a_{k+2}}\in G^{\prime}$ and the assumption that $f^{2}(a_{k})=a_{k+2}$, the element $t_{a_{k}}$ is contained in $G^{\prime}$. By $t_{a_{k}}t_{b_{k+1}}^{-1},t_{a_{k}}t_{c_{k+1}}^{-1}\in G^{\prime}$, we have $t_{b_{k+1}},t_{c_{k+1}}\in G^{\prime}$. Similarly, by $t_{b_{k+1}},t_{c_{k+1}}\in G^{\prime}$ and the assumption that $f^{-1}(b_{k+1},c_{k+1})=(b_{k},c_{k})$, we have $t_{b_{k}},t_{c_{k}}\in G^{\prime}$, and the lemma follows. ∎
Lemma 4.7.
Let $g\geq k+6$, where $k$ is a positive integer, and let $f$ be an element in $\mathcal{M}_{g}$ satisfying $f^{j}(a_{k},b_{k},c_{k})=(a_{k+j},b_{k+j},c_{k+j})$ for $j=1,2,3,4$ and $f(c_{k+4})=c_{k+5}$. We denote by $G^{\prime\prime}$ the subgroup of $\mathcal{M}_{g}$ generated by $f$ and $t_{b_{k}}t_{a_{k+2}}^{-1}t_{c_{k+3}}\cdot t_{c_{k+4}}^{-1}t_{a_{k+3}}t_{b_{k+1}}^{-1}$. Then, the elements $t_{a_{k}}t_{a_{k+1}}^{-1},t_{b_{k}}t_{b_{k+1}}^{-1},t_{c_{k}}t_{c_{k+1}}^{-1}$ are contained in $G^{\prime\prime}$.
Proof.
We set
$\Phi_{k}$ $=t_{b_{k}}t_{a_{k+2}}^{-1}t_{c_{k+3}}\cdot t_{c_{k+4}}^{-1}t_{a_{k+3}}t_{b_{k+1}}^{-1},$
$\Phi_{k+1}$ $=t_{b_{k+1}}t_{a_{k+3}}^{-1}t_{c_{k+4}}\cdot t_{c_{k+5}}^{-1}t_{a_{k+4}}t_{b_{k+2}}^{-1}.$
Note that $\Phi_{k}\in G^{\prime\prime}$. By the assumption of $f$, we see that $\Phi_{k+1}=t_{f(b_{k})}t_{f(a_{k+2})}^{-1}t_{f(c_{k+3})}\cdot t_{f(c_{k+4})}^{-1}t_{f(a_{k+3})}t_{f(b_{k+1})}^{-1}=f\Phi_{k}f^{-1}\in G^{\prime\prime}$. Since it is easy to check that
$\Phi_{k+1}\Phi_{k}(b_{k+1},a_{k+3},c_{k+4},c_{k+5},a_{k+4},b_{k+2})=(b_{k+1},a_{k+3},c_{k+4},c_{k+5},a_{k+4},a_{k+2}),$
we have
$t_{b_{k+1}}t_{a_{k+3}}^{-1}t_{c_{k+4}}\cdot t_{c_{k+5}}^{-1}t_{a_{k+4}}t_{a_{k+2}}^{-1}=(\Phi_{k+1}\Phi_{k})\Phi_{k+1}(\Phi_{k+1}\Phi_{k})^{-1}\in G^{\prime\prime}.$
Then, we obtain the following:
(1) $t_{a_{k+2}}t_{b_{k+2}}^{-1}=(t_{b_{k+1}}t_{a_{k+3}}^{-1}t_{c_{k+4}}\cdot t_{c_{k+5}}^{-1}t_{a_{k+4}}t_{a_{k+2}}^{-1})^{-1}\Phi_{k+1}\in G^{\prime\prime}.$
By $t_{a_{k+2}}t_{b_{k+2}}^{-1}\in G^{\prime\prime}$ and the assumption of $f$, we have
(2) $t_{a_{k+3}}t_{b_{k+3}}^{-1}\in G^{\prime\prime},$
(3) $t_{a_{k+4}}t_{b_{k+4}}^{-1}\in G^{\prime\prime}.$
Here, since the curves $b_{k+1},a_{k+3},c_{k+4},c_{k+5},a_{k+4}$ are disjoint from $a_{k+2}$ and $b_{k+2}$, we have $t_{b_{k+2}}^{-1}t_{a_{k+2}}=\Phi_{k+1}(t_{b_{k+1}}t_{a_{k+3}}^{-1}t_{c_{k+4}}\cdot t_{c_{k+5}}^{-1}t_{a_{k+4}}t_{a_{k+2}}^{-1})^{-1}\in G^{\prime\prime}$. By similar arguments to (2) and (3), we obtain the following:
(4) $t_{b_{k+3}}^{-1}t_{a_{k+3}}\in G^{\prime\prime},$
(5) $t_{b_{k+4}}^{-1}t_{a_{k+4}}\in G^{\prime\prime}.$
By (4), we see that $\Phi_{k}t_{a_{k+3}}^{-1}t_{b_{k+3}}=\Phi_{k}(t_{b_{k+3}}^{-1}t_{a_{k+3}})^{-1}\in G^{\prime\prime}$. Since $b_{k+1}$ is disjoint from $a_{k+3}$ and $b_{k+3}$, we have
$\Phi_{k}t_{a_{k+3}}^{-1}t_{b_{k+3}}$ $=t_{b_{k}}t_{a_{k+2}}^{-1}t_{c_{k+3}}\cdot t_{c_{k+4}}^{-1}t_{a_{k+3}}t_{b_{k+1}}^{-1}\cdot t_{a_{k+3}}^{-1}t_{b_{k+3}}$
$=t_{b_{k}}t_{a_{k+2}}^{-1}t_{c_{k+3}}\cdot t_{c_{k+4}}^{-1}t_{b_{k+3}}t_{b_{k+1}}^{-1}\ (\in G^{\prime\prime}).$
From this, we see that $(\Phi_{k}t_{a_{k+3}}^{-1}t_{b_{k+3}})^{-1}(a_{k+3},b_{k+3})=(t_{b_{k+3}}^{-1}(a_{k+3}),c_{k+3})$, and therefore, by $t_{a_{k+3}}t_{b_{k+3}}^{-1},(\Phi_{k}t_{a_{k+3}}^{-1}t_{b_{k+3}})^{-1}\in G^{\prime\prime}$, the following holds:
$t_{b_{k+3}}^{-1}t_{a_{k+3}}t_{b_{k+3}}t_{c_{k+3}}^{-1}=t_{t_{b_{k+3}}^{-1}(a_{k+3})}t_{c_{k+3}}^{-1}\in G^{\prime\prime}.$
By (4), we have
(6) $t_{b_{k+3}}t_{c_{k+3}}^{-1}\in G^{\prime\prime}.$
By (5) and $\Phi_{k},\Phi_{k+1}\in G^{\prime\prime}$, we see that $\Phi_{k}\Phi_{k+1}t_{a_{k+4}}t_{b_{k+4}}^{-1}=\Phi_{k}\Phi_{k+1}(t_{b_{k+4}}^{-1}t_{a_{k+4}})^{-1}\in G^{\prime\prime}$. Since $b_{k+2}$ are disjoint from $a_{k+4}$ and $b_{k+4}$, we have
$\Phi_{k}\Phi_{k+1}t_{a_{k+4}}t_{b_{k+4}}^{-1}$ $=t_{b_{k}}t_{a_{k+2}}^{-1}t_{c_{k+3}}\cdot t_{c_{k+5}}^{-1}t_{a_{k+4}}t_{b_{k+2}}^{-1}\cdot(t_{b_{k+4}}^{-1}t_{a_{k+4}})^{-1}$
$=t_{b_{k}}t_{a_{k+2}}^{-1}t_{c_{k+3}}\cdot t_{c_{k+5}}^{-1}t_{b_{k+4}}t_{b_{k+2}}^{-1}\ (\in G^{\prime\prime}).$
From this, we see that $(\Phi_{k}\Phi_{k+1}t_{a_{k+4}}t_{b_{k+4}}^{-1})^{-1}(a_{k+4},b_{k+4})=(t_{b_{k+4}}^{-1}(a_{k+4}),c_{k+3})$, and therefore, by $t_{a_{k+4}}t_{b_{k+4}}^{-1},(\Phi_{k}\Phi_{k+1}t_{a_{k+4}}t_{b_{k+4}}^{-1})^{-1}\in G^{\prime\prime}$, the following holds:
$t_{b_{k+4}}^{-1}t_{a_{k+4}}t_{b_{k+4}}t_{c_{k+3}}^{-1}=t_{t_{b_{k+4}}^{-1}(a_{k+4})}t_{c_{k+3}}^{-1}\in G^{\prime\prime}.$
By $t_{b_{k+4}}^{-1}t_{a_{k+4}}t_{b_{k+4}}t_{c_{k+3}}^{-1}\in G^{\prime\prime}$, (5) and the assumption that $f^{-1}(b_{k+4},c_{k+3})=(b_{k+3},c_{k+2})$, we have
(7) $t_{b_{k+4}}t_{c_{k+3}}^{-1}\in G^{\prime\prime},$
(8) $t_{b_{k+3}}t_{c_{k+2}}^{-1}\in G^{\prime\prime}.$
From (6), (7) and the assumption that $f^{-l}(b_{k+3},b_{k+4})=(b_{k+3-l},b_{k+4-l})$ for $l=1,2,3$, we obtain the following:
$t_{b_{k+3}}t_{b_{k+4}}^{-1}=t_{b_{k+3}}t_{c_{k+3}}^{-1}(t_{b_{k+4}}t_{c_{k+3}}^{-1})^{-1}\in G^{\prime\prime},$
$t_{b_{k}}t_{b_{k+1}}^{-1},t_{b_{k+1}}t_{b_{k+2}}^{-1},t_{b_{k+2}}t_{b_{k+3}}^{-1}\in G^{\prime\prime}.$
Moreover, by (1), (2), $t_{b_{k+2}}t_{b_{k+3}}^{-1}\in G^{\prime\prime}$ and the assumption that $f^{-2}(a_{k+2},a_{k+3})=(a_{k},a_{k+1})$, we have
$t_{a_{k+2}}t_{a_{k+3}}^{-1}=t_{a_{k+2}}t_{b_{k+2}}^{-1}\cdot t_{b_{k+2}}t_{b_{k+3}}^{-1}\cdot(t_{a_{k+3}}t_{b_{k+3}}^{-1})^{-1}\in G^{\prime\prime},$
$t_{a_{k}}t_{a_{k+1}}\in G^{\prime\prime}.$
By (8), (6) and the assumption that $f^{-2}(c_{k+2},c_{k+3})=(c_{k},c_{k+1})$, the following holds:
$t_{c_{k+2}}t_{c_{k+3}}^{-1}=(t_{b_{k+3}}t_{c_{k+2}}^{-1})^{-1}t_{b_{k+3}}t_{c_{k+3}}^{-1}\in G^{\prime\prime},$
$t_{c_{k}}t_{c_{k+1}}^{-1}\in G^{\prime\prime}.$
Summarizing the above, the elements $t_{a_{k}}t_{a_{k+1}}^{-1},t_{b_{k}}t_{b_{k+1}}^{-1},t_{c_{k}}t_{c_{k+1}}$ are contained in $G^{\prime\prime}$, and the lemma follows. ∎
Proof of Theorem 4.1.
Since $a_{1},b_{1},c_{1}$ are disjoint from $b_{3},a_{5},c_{6}$, we see that
$\rho_{n}(b_{3},a_{5},c_{6})=rt_{c_{1}}t_{b_{1}}^{-n}t_{a_{1}}(b_{3},a_{5},c_{6})=r(b_{3},a_{5},c_{6})=(b_{4},a_{6},c_{7}).$
This gives
$(t_{b_{3}}t_{a_{5}}^{-1}t_{c_{6}})\rho_{n}(t_{b_{3}}t_{a_{5}}^{-1}t_{c_{6}})^{-1}$ $=(t_{b_{3}}t_{a_{5}}^{-1}t_{c_{6}})\cdot\rho_{n}(t_{b_{3}}t_{a_{5}}^{-1}t_{c_{6}})^{-1}\rho_{n}^{-1}\cdot\rho_{n}$
$=(t_{b_{3}}t_{a_{5}}^{-1}t_{c_{6}})\cdot(t_{\rho_{n}(b_{3})}t_{\rho_{n}(a_{5})}^{-1}t_{\rho_{n}(c_{6})})^{-1}\cdot\rho_{n}$
$=(t_{b_{3}}t_{a_{5}}^{-1}t_{c_{6}})(t_{b_{4}}t_{a_{6}}^{-1}t_{c_{7}})^{-1}\rho_{n}.$
and therefore, by $\rho_{n},(t_{b_{3}}t_{a_{5}}^{-1}t_{c_{6}})\rho_{n}(t_{b_{3}}t_{a_{5}}^{-1}t_{c_{6}})^{-1}\in G_{1}$, $G_{1}$ contains the element $(t_{b_{3}}t_{a_{5}}^{-1}t_{c_{6}})(t_{b_{4}}t_{a_{6}}^{-1}t_{c_{7}})^{-1}$. From Lemma 4.7, we have $t_{a_{3}}t_{a_{4}}^{-1},t_{b_{3}}t_{b_{4}}^{-1},t_{c_{3}}t_{c_{4}}^{-1}\in G_{1}$ (by letting $k=3$ and $f=\rho_{n}$). Moreover, from Lemma 4.6, we see that the elements $t_{a_{3}},t_{b_{3}},t_{c_{3}}$ are in $G_{1}$ (by letting $k=3$ and $f=\rho_{n}$). Since it is easy to check that $\rho_{n}(a_{i},b_{i})=(a_{i+1},b_{i+1})$ for $i=3,4,\ldots,g-1$ and $\rho_{n}(a_{g},b_{g})=(a_{1},b_{1})$, by $t_{a_{3}},t_{b_{3}},\rho_{n}^{g-2}\in G_{1}$, we have
$t_{a_{1}}\in G_{1},$
$t_{b_{1}}\in G_{1}.$
Hence, we obtain $rt_{c_{1}}=rt_{c_{1}}t_{b_{1}}^{-n}t_{a_{1}}\cdot t_{a_{1}}^{-1}t_{b_{1}}^{n}=\rho_{n}t_{a_{1}}^{-1}t_{b_{1}}^{n}\in G_{1}$. By $(rt_{c_{1}})^{-2}(c_{3})=c_{1}$, we have
$t_{c_{1}}$ $=t_{(rt_{c_{1}})^{-2}(c_{3})}=(rt_{c_{1}})^{-2}t_{c_{3}}(rt_{c_{1}})^{2}\in G_{1},$
$r$ $=rt_{c_{1}}\cdot t_{c_{1}}^{-1}\in G_{1}.$
Since $r^{i}(a_{1},b_{1},c_{1})=(a_{i+1},b_{i+1},c_{i+1})$ for $i=0,1,2,\ldots,g-1$, we have
$t_{a_{i+1}},t_{b_{i+1}},t_{c_{i+1}}\in G_{1}$
for $i=0,1,2,\ldots,g-1$. Therefore, we see that $G_{1}=\mathcal{M}_{g}$, which proves the theorem. ∎
Proof of Theorem 4.3.
We omit the proof since it is almost the same as the proof of Theorem 4.1. ∎
Proof of Theorem 4.2.
Since $a_{1},b_{1},c_{1}$ are disjoint from $a_{2},c_{2},a_{3},b_{3}$, we see that $\rho_{n}(a_{2},c_{2},a_{3},b_{3})=rt_{c_{1}}t_{b_{1}}^{-n}t_{a_{1}}(a_{2},a_{3},b_{3},c_{2})=r(a_{2},a_{3},b_{3},c_{2})$, and hence
$\rho_{n}(a_{2})$ $=a_{3},$
$\rho_{n}(a_{3},b_{3},c_{2})$ =$=\begin{cases}(a_{4},b_{4},c_{3})&(g\geq 4),\\ (a_{1},b_{1},c_{3})&(g=3).\end{cases}$
Since we have
$t_{a_{2}}\rho_{n}t_{a_{2}}^{-1}$ $=t_{a_{2}}(\rho_{n}t_{a_{2}}^{-1}\rho_{n}^{-1})\rho_{n}$
$(t_{a_{3}}t_{b_{3}}t_{c_{2}})\rho_{n}(t_{a_{3}}t_{b_{3}}t_{c_{2}})^{-1}$ $=(t_{a_{3}}t_{b_{3}}t_{c_{2}})\{\rho_{n}(t_{a_{3}}t_{b_{3}}t_{c_{2}})^{-1}\rho_{n}^{-1}\}\rho_{n},$
from the above arguments and the relation $t_{f(c)}=ft_{c}f^{-1}$, the following holds:
$G_{2}\ni t_{a_{2}}\rho_{n}t_{a_{2}}^{-1}=t_{a_{2}}t_{a_{3}}^{-1}\rho_{n},$
$G_{2}\ni(t_{a_{3}}t_{b_{3}}t_{c_{2}})\rho_{n}(t_{a_{3}}t_{b_{3}}t_{c_{2}})^{-1}=\begin{cases}(t_{a_{3}}t_{b_{3}}t_{c_{2}})\left(t_{a_{4}}t_{b_{4}}t_{c_{3}}\right)^{-1}\rho_{n}&(g\geq 4),\\ (t_{a_{3}}t_{b_{3}}t_{c_{2}})\left(t_{a_{1}}t_{b_{1}}t_{c_{3}}\right)^{-1}\rho_{n}&(g=3).\end{cases}$
Therefore, we have
(9) $t_{a_{2}}t_{a_{3}}^{-1}\in G_{2},$
$t_{a_{3}}t_{b_{3}}t_{c_{2}}\cdot t_{c_{3}}^{-1}t_{b_{4}}^{-1}t_{a_{4}}^{-1}\in G_{2}\ \mathrm{if}\ g\geq 4,$
$t_{a_{3}}t_{b_{3}}t_{c_{2}}\cdot t_{c_{3}}^{-1}t_{b_{1}}^{-1}t_{a_{1}}^{-1}\in G_{2}\ \mathrm{if}\ g=3.$
Hereafter, we prove that $t_{a_{3}}\in G_{2}$ using these elements and Lemma 4.5. More precisely, we show that
$t_{a_{1}}t_{a_{2}}^{-1},\ t_{a_{1}}t_{a_{3}}^{-1},\ t_{a_{1}}t_{b_{2}}^{-1},\ t_{a_{1}}t_{c_{1}}^{-1},\ t_{a_{2}}t_{c_{2}}^{-1},\ t_{a_{1}}t_{c_{2}}^{-1},\ t_{a_{1}}t_{b_{3}}^{-1}\in G_{2}.$
Suppose that $g\geq 4$. For simplicity of notation, we write $\phi=t_{a_{3}}t_{b_{3}}t_{c_{2}}\cdot t_{c_{3}}^{-1}t_{b_{4}}^{-1}t_{a_{4}}^{-1}(\in G_{2})$.
We first show that $t_{a_{1}}t_{a_{2}}^{-1},t_{a_{1}}t_{a_{3}}^{-1},t_{a_{1}}t_{b_{3}}^{-1}\in G_{2}$. It is easy to check that $\phi(a_{2},a_{3})=(a_{2},b_{3})$, and hence we have $t_{a_{2}}t_{b_{3}}^{-1}\in G_{2}$ by $t_{a_{2}}t_{a_{3}}^{-1},\phi\in G_{2}$. Since $a_{1},b_{1},c_{1}$ are disjoint from $a_{i_{1}}$ and $b_{i_{2}}$ for $i_{1}=2,3,\ldots,g$ and $i_{2}=3,4,\ldots,g$, we see that $\rho_{n}(a_{i},a_{i+1},b_{i+1})=rt_{c_{1}}t_{b_{1}}^{-n}t_{a_{1}}(a_{i},a_{i+1},b_{i+1})=(a_{i+1},a_{i+2},b_{i+2})$ for $i=2,3,\ldots,g-2$ and $\rho_{n}(a_{g},b_{g})=rt_{c_{1}}t_{b_{1}}^{-n}t_{a_{1}}(a_{g},b_{g})=(a_{1},b_{1})$. From the above arguments and $\rho_{n}^{i},t_{a_{2}}t_{a_{3}}^{-1},t_{a_{2}}t_{b_{3}}^{-1}\in G_{2}$ for $i=1,2,\ldots,g-2$, we have
(10) $t_{a_{2+i}}t_{a_{3+i}}^{-1}\in G_{2}\ \ (0\leq i\leq g-3)$ $\mathrm{and}$ $t_{a_{g}}t_{a_{1}}^{-1}\in G_{2}\ (i=g-2),$
(11) $t_{a_{g}}t_{b_{1}}^{-1}\in G_{2}\ (i=g-2).$
Using these, we have
$t_{a_{2}}t_{a_{1}}^{-1}=t_{a_{2}}t_{a_{3}}^{-1}\cdot t_{a_{3}}t_{a_{4}}^{-1}\cdots t_{a_{g-1}}t_{a_{g}}^{-1}\cdot t_{a_{g}}t_{a_{1}}^{-1}\in G_{2}.$
Since $a_{1}$ (resp. $a_{g}$) is disjoint from $a_{2}$ (resp. $a_{1}$ and $b_{1}$), the following holds:
(12) $t_{a_{1}}t_{a_{2}}^{-1}=t_{a_{2}}t_{a_{1}}^{-1}\in G_{2},$
(13) $t_{a_{1}}t_{a_{3}}^{-1}=t_{a_{1}}t_{a_{2}}^{-1}\cdot t_{a_{2}}t_{a_{3}}^{-1}\in G_{2},$
(14) $t_{a_{1}}^{-1}t_{b_{1}}=(t_{a_{g}}t_{a_{1}}^{-1})\cdot(t_{a_{g}}t_{b_{1}}^{-1})^{-1}\in G_{2},$
$t_{a_{1}}t_{b_{1}}^{-1}=(t_{a_{g}}t_{a_{1}}^{-1})^{-1}(t_{a_{g}}t_{b_{1}}^{-1})\in G_{2}.$
Moreover, since $\rho_{n}^{-g+2}(a_{1},b_{1})=(a_{3},b_{3})$, we have
(15) $t_{a_{3}}t_{b_{3}}^{-1}=t_{\rho_{n}^{-g+2}(a_{1})}t_{\rho_{n}^{-g+2}(b_{1})}^{-1}=\rho_{n}^{-g+2}t_{a_{1}}t_{b_{1}}^{-1}\rho_{n}^{g-2}\in G_{2}.$
Therefore, we see that
(16) $t_{a_{1}}t_{b_{3}}^{-1}=t_{a_{1}}t_{a_{3}}^{-1}\cdot t_{a_{3}}t_{b_{3}}^{-1}\in G_{2}.$
From (12), (13) and (16), we have $t_{a_{1}}t_{a_{2}}^{-1},t_{a_{1}}t_{a_{3}}^{-1},t_{a_{1}}t_{b_{3}}^{-1}\in G_{2}$.
Next, we show that $t_{a_{1}}t_{b_{2}}^{-1},t_{a_{1}}t_{c_{1}}^{-1},t_{a_{2}}t_{c_{2}}^{-1},t_{a_{1}}t_{c_{2}}^{-1}\in G_{2}$. By (9) and (15), we have $t_{a_{2}}t_{b_{3}}^{-1}=t_{a_{2}}t_{a_{3}}^{-1}\cdot t_{a_{3}}t_{b_{3}}^{-1}\in G_{2}$, and hence, by $\rho_{n}(a_{2},b_{3})=(a_{3},b_{4})$ and $t_{a_{2}}t_{b_{3}}^{-1},\rho_{n}\in G_{2}$, we have $t_{a_{3}}t_{b_{4}}^{-1}\in G_{2}$. Combining this and (13), the element $t_{a_{1}}t_{b_{4}}^{-1}=t_{a_{1}}t_{a_{3}}^{-1}\cdot t_{a_{3}}t_{b_{4}}^{-1}$ is in $G_{2}$. Since it is easily seen that $\phi^{-1}(a_{1},b_{4})=(a_{1},c_{3})$, we have $t_{a_{1}}t_{c_{3}}^{-1}\in G_{2}$ by $t_{a_{1}}t_{b_{4}}^{-1},\phi^{-1}\in G_{2}$. Using this, (12) and (13), we have
$t_{a_{2}}t_{c_{3}}^{-1}=(t_{a_{1}}t_{a_{2}}^{-1})^{-1}t_{a_{1}}t_{c_{3}}^{-1}\in G_{2},$
$t_{a_{3}}t_{c_{3}}^{-1}=(t_{a_{1}}t_{a_{3}}^{-1})^{-1}t_{a_{1}}t_{c_{3}}^{-1}\in G_{2}.$
Here, we set $\sigma_{n}=\rho_{n}\cdot t_{a_{1}}^{-1}t_{b_{1}}\cdot(t_{a_{2}}^{-1}t_{a_{1}}\cdot t_{a_{1}}^{-1}t_{b_{1}})^{n-1}$. Then, by the disjointness of $a_{2}$ and $b_{1}$, (12) and (14), we have $G_{2}\ni\sigma_{n}=rt_{c_{1}}t_{a_{2}}^{-n+1}$. Since it is obvious that $\sigma_{n}^{-2}(a_{3},c_{3})=(a_{1},c_{1})$, $\sigma_{n}^{-1}(a_{3},c_{3})=(a_{2},c_{2})$ and $\sigma_{n}^{-2}(a_{2},c_{3})=(a_{1},c_{2})$, by $t_{a_{3}}t_{c_{3}}^{-1},t_{a_{2}}t_{c_{3}}^{-1},\sigma_{n}\in G_{2}$, we see that the following holds:
(17) $t_{a_{1}}t_{c_{1}}^{-1}\in G_{2},$
(18) $t_{a_{2}}t_{c_{2}}^{-1}\in G_{2},$
(19) $t_{a_{1}}t_{c_{2}}^{-1}\in G_{2}.$
Moreover, since it is immediate that $\sigma_{n}(a_{1},b_{1})=(a_{2},t_{c_{2}}(b_{2}))$, by (14) and $\sigma_{n}\in G_{2}$ we have $t_{a_{2}}^{-1}t_{t_{c_{2}}(b_{2})}\in G_{2}$. Since $t_{a_{2}}^{-1}t_{t_{c_{2}}(b_{2})}=t_{a_{2}}^{-1}t_{c_{2}}t_{b_{2}}t_{c_{2}}^{-1}$ by the relation $t_{t_{c_{2}}(b_{2})}=t_{c_{2}}t_{b_{2}}t_{c_{2}}^{-1}$, we have $t_{c_{2}}t_{b_{2}}^{-1}\in G_{2}$ by (18) and the disjointness of $a_{2}$ and $c_{2}$. Therefore, by (19) the element $t_{a_{1}}t_{b_{2}}^{-1}=t_{a_{1}}t_{c_{2}}^{-1}\cdot t_{c_{2}}t_{b_{2}}^{-1}$ is contained in $G_{2}$. Summarizing the above arguments, we have $t_{a_{1}}t_{b_{2}}^{-1},t_{a_{1}}t_{c_{1}}^{-1},t_{a_{2}}t_{c_{2}}^{-1},t_{a_{1}}t_{c_{2}}^{-1}\in G_{2}$.
By Lemma 4.5 we see that $t_{a_{3}}\in G_{2}$. Moreover, by (13), (14) and (17) we have
$t_{a_{1}}$ $=t_{a_{1}}t_{a_{3}}^{-1}\cdot t_{a_{3}}\in G_{2},$
$t_{b_{1}}$ $=t_{a_{1}}\cdot t_{a_{1}}^{-1}t_{b_{1}}\in G_{2},$
$t_{c_{1}}$ $=(t_{a_{1}}t_{c_{1}}^{-1})^{-1}\cdot t_{a_{1}}^{-1}\in G_{2},$
and hence $r=\rho_{n}t_{a_{1}}^{-1}t_{b_{1}}^{n}t_{c_{1}}^{-1}\in G_{2}$. Since $r^{i}(a_{1},b_{1},c_{1})=(a_{i+1},b_{i+1},c_{i+1})$ for $i=0,1,2,\ldots,g-1$, we have
$t_{a_{i+1}},t_{b_{i+1}},t_{c_{i+1}}\in G_{2}$
for $i=0,1,2,\ldots,g-1$. Therefore, we see that $G_{2}=\mathcal{M}_{g}$ for $g\geq 4$.
Suppose that $g=3$. For simplicity of notation, we also write $\phi=t_{a_{3}}t_{b_{3}}t_{c_{2}}\cdot t_{c_{3}}^{-1}t_{b_{1}}^{-1}t_{a_{1}}^{-1}(\in G_{2})$. Since it is easy to check that $\rho_{n}(a_{2},a_{3})=(a_{3},a_{1})$, by (9) and $\rho_{n}\in G_{2}$ we have
(20) $t_{a_{3}}t_{a_{1}}^{-1}\in G_{2},$
and hence by (9), $\phi(a_{2},a_{3})=(a_{2},b_{3})$ and $\phi\in G_{2}$, we see that $t_{a_{2}}t_{b_{3}}^{-1}\in G_{2}$. From this, $\rho_{n}(a_{2},b_{3})=(a_{3},b_{1})$ and $\rho_{n}\in G_{2}$, we have
(21) $t_{a_{3}}t_{b_{1}}^{-1}\in G_{2}.$
Moreover, the following holds:
(22) $t_{a_{1}}t_{a_{3}}^{-1}=(t_{a_{3}}t_{a_{1}}^{-1})^{-1}\in G_{2},$
(23) $t_{a_{1}}t_{a_{2}}^{-1}=t_{a_{1}}t_{a_{3}}^{-1}\cdot(t_{a_{2}}t_{a_{3}}^{-1})^{-1}\in G_{2},$
(24) $t_{a_{1}}t_{b_{3}}^{-1}=t_{a_{1}}t_{a_{2}}^{-1}\cdot t_{a_{2}}t_{b_{3}}^{-1}\in G_{2}.$
By (9), (20), (21) and the disjointness of $a_{3}$ and $a_{1},b_{1}$, we have
(25) $t_{a_{1}}^{-1}t_{b_{1}}=t_{a_{3}}t_{a_{1}}^{-1}\cdot(t_{a_{3}}t_{b_{1}}^{-1})^{-1}\in G_{2},$
(26) $t_{a_{2}}t_{b_{1}}^{-1}=t_{a_{2}}t_{a_{3}}^{-1}\cdot t_{a_{3}}t_{b_{1}}^{-1}\in G_{2}$
Here, when we set $\sigma_{n}=\rho_{n}\cdot t_{a_{1}}^{-1}t_{b_{1}}(t_{a_{2}}t_{b_{1}}^{-1})^{-n+1}$, we see that $\sigma_{n}=rt_{c_{1}}t_{a_{2}}^{-n+1}\in G_{2}$ by (26) and the disjointness of $a_{2}$ and $b_{1}$. Since it is obvious that $\phi^{-1}(a_{2},b_{1})=(a_{2},c_{3})$, by (26) and $\phi^{-1}\in G_{2}$ we have $t_{a_{2}}t_{c_{3}}^{-1}\in G_{2}$. Moreover, by $\sigma_{n}(a_{2},c_{3})=(a_{3},c_{1}),\sigma_{n}(a_{3},c_{1})=(a_{1},c_{2})$ and $t_{a_{2}}t_{c_{3}}^{-1},\sigma_{n}\in G_{2}$ we see that
$t_{a_{3}}t_{c_{1}}^{-1}\in G_{2},$
(27) $t_{a_{1}}t_{c_{2}}^{-1}\in G_{2}.$
From these, (22) and (23), we have
(28) $t_{a_{1}}t_{c_{1}}^{-1}=t_{a_{1}}t_{a_{3}}^{-1}\cdot t_{a_{3}}t_{c_{1}}^{-1}\in G_{2},$
(29) $t_{a_{2}}t_{c_{2}}^{-1}=(t_{a_{1}}t_{a_{2}}^{-1})^{-1}\cdot t_{a_{1}}t_{c_{2}}^{-1}\in G_{2}.$
Since it is immediate that $\sigma_{n}(a_{1},b_{1})=(a_{2},t_{c_{2}}(b_{2}))$, by (25) and $\sigma_{n}\in G_{2}$ we have
$G_{2}\ni t_{a_{2}}^{-1}t_{t_{c_{2}}(b_{2})}=t_{a_{2}}^{-1}t_{c_{2}}t_{b_{2}}t_{c_{2}}^{-1}=(t_{c_{2}}^{-1}t_{a_{2}})^{-1}t_{b_{2}}t_{c_{2}}^{-1}.$
By (29) and the disjointness of $a_{2}$ and $c_{2}$, we see that $t_{b_{2}}t_{c_{2}}^{-1}\in G_{2}$. Therefore, by (27), the following holds:
$t_{a_{1}}t_{b_{2}}^{-1}=t_{a_{1}}t_{c_{2}}^{-1}\cdot(t_{b_{2}}t_{c_{2}}^{-1})^{-1}\in G_{2}.$
From (22)–(24), (27)–(29) and this, we have $t_{a_{3}}\in G_{2}$ using Lemma 4.5. By a similar argument to that in the case of $g\geq 4$, we see that $G_{2}=\mathcal{M}_{3}$. ∎
Proof of Theorem 4.4.
Suppose that $g\geq 4$. By a similar argument to the proof of Theorem 4.2, we see that $t_{a_{1}}t_{a_{2}}^{-1},t_{a_{1}}t_{a_{3}}^{-1},t_{a_{1}}t_{b_{3}}^{-1}\in G_{4}$. Moreover, the element $\rho^{\prime}=rt_{c_{1}}$ is equal to $\sigma_{1}$ in the proof of Theorem 4.2. Therefore, we have $t_{a_{1}}t_{b_{2}}^{-1},$ $t_{a_{1}}t_{c_{1}}^{-1},$ $t_{a_{2}}t_{c_{2}}^{-1},$ $t_{a_{1}}t_{c_{2}}^{-1}\in G_{4}$, and hence the elements $t_{a_{1}},t_{b_{1}},t_{c_{1}},r$ are contained in $G_{4}$. By a similar argument to the proof of Theorem 4.1, we see that $G_{4}=\mathcal{M}_{g}$ for $g\geq 4$. The same reasoning applies to the case $g=3$. This finishes the proof. ∎